### v1

# ToxicTags: Decoding Toxic Memes with Rich Tag Annotations

Abstract
The 2025 Global Risks Report [^1] https://reports.weforum.org/docs/WEF˙Global˙Risks˙Report˙2025.pdf identifies state-based armed conflict and societal polarisation among the most pressing global threats, with social media playing a central role in amplifying toxic discourse. Memes, as a widely used mode of online communication, often serve as vehicles for spreading harmful content. However, limitations in data accessibility and the high cost of dataset curation hinder the development of robust meme moderation systems. To address this challenge, in this work, we introduce a first-of-its-kind dataset of 6,300 real-world meme-based posts annotated in two stages: (i) binary classification into toxic and normal, and (ii) fine-grained labelling of toxic memes as hateful, dangerous, or offensive. A key feature of this dataset is that it is enriched with auxiliary metadata of socially relevant tags, enhancing the context of each meme. In addition, we propose a tag generation module that produces socially grounded tags, because most in-the-wild memes often do not come with tags. Experimental results show that incorporating these tags substantially enhances the performance of state-of-the-art VLMs in toxicity detection tasks. Our contributions offer a novel and scalable foundation for improved content moderation in multimodal online environments. Warning: Contains potentially toxic contents.
1 Introduction
While communication on online platforms spans a variety of modalities [^2] https://www.sprinklr.com/blog/types-of-social-media/, memes have emerged as a popular and influential form of expression – initially intended for lighthearted humor. However, they are increasingly being misused as vehicles for spreading harmful content, including hate speech, misinformation, and toxic ideologies.
Identifying such harmful content is particularly challenging due to the subtle and context-dependent nature of memes. Their meaning is often embedded in cultural references, online trends, sarcasm, or coded language, making them difficult to interpret not only for automated systems but even for human moderators. In response to these challenges, vision language models (VLMs) have recently gained traction as powerful tools for content moderation. These models are capable of jointly analyzing visual and textual elements to grasp the nuanced context of memes and can provide detailed justifications for their classifications (Qu, Backes, and Zhang 2025). Despite these advances, recent research highlights the limitations of VLMs in accurately detecting hateful memes (Rizwan et al. 2025), underscoring the urgent need to bridge these performance gaps. Similarly, a recent blog post [^3] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes/ by Meta highlighted the challenges and complexities inherent in their current content moderation frameworks.
Dataset scarcity: Datasets serve as the foundational fuel for generative AI models. However, researchers increasingly face significant obstacles in curating such datasets, primarily due to the high costs of manual annotation and the limitations imposed on large-scale crawling of social media platforms. These challenges have resulted in datasets that are either manually constructed (Kiela et al. 2020; Lin et al. 2024) – often failing to fully capture the richness of human creativity – or narrowly scoped to specific targets or events (Pramanick et al. 2021a, b; Fersini et al. 2022). Moreover, this scarcity of comprehensive datasets has hindered efforts to simulate the complexity of social media ecosystems across a broad taxonomy of content labels. This is in contrast to textual hate speech research, where more extensive taxonomical explorations exist (Sachdeva et al. 2022).
Our contributions: To address the aforementioned limitations, we make the following contributions.
We introduce a novel, highly diverse, real-world meme dataset enriched with human-annotated contextual tags – a critical yet often overlooked feature in social media content. Unlike prior datasets, it consists of only real-world memes with no restrictions based on specific targets or events (see Section 3). The final dataset comprises 6,300 annotated memes, capturing a wide spectrum of online discourse.
We design a rigorous two-stage human annotation pipeline. In the first stage, memes are classified as either toxic or normal. In the second stage, toxic memes are further categorized into one of three fine-grained classes: hateful, dangerous, or offensive. These categories have been distilled through an iterative annotation process, revealing that a four-class taxonomy is appropriately expressive for moderating a wide range of social media content. This taxonomy can help reduce misclassifications in existing moderation pipelines.
We propose a novel tag generation framework (see Section 6) significantly departing from existing approaches. Leveraging real-world context, our system integrates Google Search API and Google Lens to produce high-quality, socially grounded tags for memes.
We perform few-shot prompting of VLMs using various schemes to select the samples. Our experiments demonstrate that the inclusion of contextual tags significantly enhances model performance in meme classification tasks.
Finally, we report the performance of our method on a standard hate meme detection dataset. Note that this dataset does not have the tag information in the first place; therefore, we use our tag generation module to predict relevant tags for a query image. Inclusion of these tags in the scheme for sampling the few-shot examples results in a better performance compared to the case where the tags are absent.
2 Related works
Hate meme detection: Rapid surge of hate around the globe (Arcila Calderón et al. 2024) in the recent past years, with memes acting as a major source of fuel, has led to the curation of multiple hateful memes dataset (Kiela et al. 2020; Lin et al. 2025). Similarly, there has been extensive research in the field to build robust content moderation frameworks (Rizwan et al. 2025; Das and Mukherjee 2023b; Prakash et al. 2023; Cao, Lee, and Jiang 2024; Cao et al. 2023), with some works on low-resource languages (Das and Mukherjee 2023a; Kumari et al. 2024). Despite such rapid developments, current datasets are either limited to manually curated memes or focus only on a subset of events (Pramanick et al. 2021a, b; Chen et al. 2023).
dataset
event/target
independent?
real-world
memes?
post’s contextual
information?
two stage
annotation?
size labels
FHM (Kiela et al. 2020) $\sim$10,000 hateful, not-hateful
MAMI (Fersini et al. 2022) $\sim$10,000 misogynistic, not-misogynistic
HARM-C (Pramanick et al. 2021a) $\sim$3,544 very harmful, partially harmful, harmless
HARM-P (Pramanick et al. 2021b) $\sim$3,552 very harmful, partially harmful, harmless
UA-RU Conflict  (Thapa et al. 2022) $\sim$5,680 hateful, not-hateful
CrisisHateMM (Bhandari et al. 2023) $\sim$4,723 hateful, not-hateful
RUHate-MM (Thapa et al. 2024) $\sim$20,675 hateful, not-hateful
ToxicTags (ours) $\sim$6,300
I– toxic, normal
II– hateful, dangerous, offensive, normal
Table 1: Comparison of our dataset with existing datasets on toxic memes. Here post’s contextual information refers to the title or tags associated with the meme (as in ToxicTags).
Tags: Prior research has extensively explored automated tag generation for images (Huang et al. 2024; Zhang et al. 2024; Dai et al. 2023), leading to the development of several dedicated models and datasets. However, to the best of our knowledge, no existing work addresses the specific challenge of tag generation for memes – visual artefacts that often combine text and imagery in culturally nuanced, context-dependent ways. To fill this gap, we curate a dedicated meme-tagging dataset and introduce a novel framework built around KeyLLM [^4] https://maartengr.github.io/KeyBERT/guides/keyllm.html to generate socially relevant and context-aware tags from memes (see Section 6).
This work: We present, for the first time, a diverse real-world dataset of memes labelled across four nuanced categories – toxic, hateful, dangerous, offensive – alongside normal, moving beyond the traditional binary classification used in prior studies (refer to Table 1). In addition, we introduce a novel, socially aware tag generation module, opening new avenues in open-set image tagging. Finally, we employ these tags to improve the performance of toxic meme detection.
3 Dataset
Collection: We source real-world social media memes from [https://imgflip.com], specifically utilizing its streams section [^5] https://imgflip.com/streams to crawl meme-centric posts. The platform was chosen for two primary reasons: (i) its strong emphasis on social media-style interactions where memes serve as a central mode of communication, and (ii) its wide variety of user-generated content, allowing us to collect memes without any target-specific or event-specific filtering, thereby closely mirroring real-world social media environments (refer to Table 1 for comparison with existing datasets).
For our dataset, we select three high-volume and thematically rich streams [^6] https://imgflip.com/m/Dark˙humour, https://imgflip.com/m/MEMES˙OVERLOAD, https://imgflip.com/m/politicsDark_Humour, Memes_Overload, and Politics based on their popularity and relevance to diverse meme content. This selection has been made through manual inspection of stream descriptions and content samples (see Appendix C for details). Further, to ensure that the memes included are socially engaging and contextually rich, we manually review a subset of posts and observe that memes with fewer than two comments are often unengaging or irrelevant to a broader audience. Hence, we retain only those memes that received at least two comments. In total, we manually curate 37,072 memes across the selected streams over a period of approximately one month. The memes were downloaded using a browser-based image downloader extension [^7] https://chromewebstore.google.com/detail/download-all-images/ifipmflagepipjokmbdecpmjbibjnakm?hl=en. As a result, our dataset offers a large and diverse collection of real-world memes, free from artificial filtering or event-driven bias.
Metadata: For each post, along with the meme, we collect its (i) title, (ii) number of comments and (iii) the tags list. For the collected memes, we extract the embedded text using GoogleOCR [^8] https://cloud.google.com/vision/docs/ocr. To ensure robustness of OCR extracted text, we perform manual verification of 100 randomly chosen samples and find nearly 98 to be correctly identified; thus depicting the robustness of the tool. Further, we also store the bounding box of OCR and use it later for generating tags (refer to Section 6).
Pre-processing: Before providing the meme along with its metadata for annotation, we perform the following pre-processing steps.
(i) Deduplication: Out of the initially collected 37,072 posts, we first apply exact string matching on the title and tags fields to identify duplicate entries. Subsequently, we perform visual deduplication on the matched posts using the imagededup [^9] https://idealo.github.io/imagededup/methods/hashing/ library, employing a Hamming distance threshold of zero to ensure strict removal of visually identical images. This two-step deduplication process – textual followed by visual – ensures that only unique memes, along with their associated metadata, are retained. After this filtering, we obtain a final dataset comprising 6,300 unique and high-quality meme samples.
(ii) Removal of unwanted tags: To reduce noise and enhance the quality of our tags, commonly occurring irrelevant tags like – ‘darkhumour’, ‘memes’, ‘you have been eternally cursed for reading the tags’ – are manually removed from each post by expert researchers (refer to Section 4). Subsequently, we obtain a rich set of 7,209 unique and socially relevant tags from an initial list of 9,664 unique tags.
Agreement: We strictly adhere to the privacy and copyright regulations of the platform [^10] https://imgflip.com/terms and therefore collected our data only from the publicly available posts. To maintain the privacy of users, we did not store any information that could potentially compromise their anonymity.
4 Annotation
Two staged annotation: We adopt a two stage annotation process – (i) classification of each meme into toxic or normal, and (ii) further bifurcation of toxic memes into offensive, dangerous or hateful. The two-stage process is specifically adopted to mitigate annotation errors as suggested in (Rizwan et al. 2025). Further, we also conduct a pilot study at each stage before concluding the final annotations. For both stages, we perform three annotations per sample by three different annotators to minimise annotation bias, given that this is a highly subjective task (Mathew et al. 2022). We use the standard definition of labels that are effectively used for practical applications. In the upcoming subsections, we discuss each stage separately and further present detailed annotation instructions with corresponding definitions of the labels in Appendix D (also see Figure 5 in the Appendix).
Annotator details: We initially selected 25 annotators through an open call in a university setting. The minimum criteria that we set are (i) the enrollment of the candidate in the second year of their bachelor’s program, and (ii) familiarity of the candidate with diverse social media content. Among these, five are experienced researchers having exposure to this field, and the rest 20 candidates are engineering students in their sophomore years. All the annotators are within the age range of 21-27 years, and the whole annotation process has been conducted under the close supervision of two senior researchers. To help the annotators with their regular queries and to make sure their mental health is not affected, we conducted daily scrum and further made a Slack workspace for easier collaboration. In particular, we asked all the participating annotators to sign an agreement before starting the annotation, wherein we specifically focus on their mental well-being, given the nature of the work. We detail out the safety guidelines in Appendix D; further, as an extra safety measure, we rolled out a maximum of only 50 samples per day to each annotator.
Annotation tool: Separate PDFs for memes with all required metadata (as shown in Figure 5 in the Appendix) and a separate Google sheet have been provided to each annotator to reduce bias and ensure fair annotation. Further in our scrum with the annotators, we assisted them by discussing their queries to provide them further insights so as to maintain a high quality of annotation.
Annotation codebook: We also provide detailed annotation instructions, accompanied by 25 samples that have been manually selected and annotated by the two supervising researchers. As outlined in Appendix D, our annotation codebook has been constructed based on standard guidelines from which we have derived our definitions. For the toxic label, we strictly follow the definition used by the Perspective API, given its widespread adoption in production settings. For the hateful label, we adhere to the Facebook hate meme definition (Kiela et al. 2020), which has become the de facto standard. For the dangerous label, we follow the annotation guidelines available at the following source [^11] https://www.dangerousspeech.org/dangerous-speech.
Stage I: Binary labelling of the dataset
Annotation: The full set of 6,300 samples has been evenly distributed among the 12 selected annotators (out of the 25 recruited). Each post has been independently annotated by three annotators. We achieve a final inter-annotator agreement score of 0.8176, as measured by Fleiss’ $\kappa$. Notably, for a subjective task (Mathew et al. 2022) such as ours, this level of agreement is considered relatively high, thereby validating the quality of our annotations. Each annotator has been paid USD 40 for this task which is much above the minimum wage in the annotators’ country.
Label assignment: The final label for each sample is determined based on the majority vote of the three annotations, i.e., the label with at least two agreements is considered to be the final label. In this stage, we finally obtained 1,392 normal and 4908 toxic samples (refer to Table 2 for complete statistics).
Stage II: Fine-grained labelling
Annotation: As a pilot step, we randomly sampled 250 toxic memes from the previous stage and had the expert annotators categorise them into three distinct labels. These labelled samples were then evenly distributed among the 12 annotators, each of whom annotated all the assigned samples. Upon verification, we found that the annotators were performing satisfactorily, and thus proceeded with the final annotation with this group. Subsequently, the set of 4,908 toxic samples was categorised into three classes: hateful, dangerous, and offensive. As before, each sample received three independent annotations. The annotators achieved a Fleiss’ $\kappa$ agreement score of 0.8197 – a slight improvement over Stage I – highlighting the value of multi-stage and multi-pilot studies in enhancing annotation quality.
Label assignment: The final label is determined based on the majority vote of the annotators for each sample. When no label receives more than one vote (which means equal vote to hateful, dangerous and offensive for that particular sample), we exclude that sample from the dataset and put it in undecided category. The overall dataset statistics are noted in Table 2. To maintain the diversity of tags in both the splits, we ensure that each tag appears at least 15% of the times in the test split, resulting in a total of 1000 test samples and 5,300 training samples. Word clouds for label wise tags are presented in Table 9 of Appendix E.
stage label train test total
I & II normal 1243 149 1392
I toxic 4057 851 4908
hateful 1475 343 1818
dangerous 1788 375 2163
offensive 653 122 775
II undecided 141 11 152
Total 5300 1000 6300
Table 2: Data statistics showing the distribution of the data points across the classes.
5 Analysis of the dataset
Figure 1: Frequently occurring tags in the toxic, hateful, offensive, and normal classes.
One of the most unique features of our dataset is the tags associated with each meme. This allows us to perform certain interesting analysis that we report below. Some of the most frequent tags we find in the train and the test splits are illustrated in Figure 1. We report the top 30 tags from the toxicity class as well as the top 10 tags each from the hateful, dangerous, offensive and normal classes, respectively. Some properties of the top tags are enlisted below (see Table 11 in the Appendix for more details).
(i) The distribution of the top tags across the train and the test splits is very similar, indicating that our strategy of splitting the data is effective.
(ii) The most prevalent tags in the hateful class are ‘9/11’, ‘nazi’, ‘adolf hitler’, ‘pedophile’ and ‘racist’ indicating their popularity in spreading hateful sentiments. We also observe that for the dangerous class, some tags that appear benign at the surface level and are primarily associated with children’s media show up. These include ‘ernie and bert’, ‘sesame street’, and ‘sonic the hedgehog’ which are used to render a comic angle to the dangerous posts (see Table 10 in the Appendix for more examples).”
(iii) We observe that users are generally reluctant to reshare content involving serious violence, suicide, or murder, as indicated by the relatively low frequency of the ‘repost’ tag in the dangerous category compared to its prevalence in other categories.
(iv) The widespread use of the ‘lol’ tag suggests an attempt to downplay harmful content through humour.
(v) Frequent use of tags such as ‘cannibalism’, ‘murder’, ‘suicide’, ‘abortion’, and violent phrases like ‘school shooting’ underscores how dark humor is often employed as a tool to normalize or propagate digital violence, thereby contributing to a heightened sense of danger.
(vi) One notable observation concerns the use of the tag ‘alabama’, which frequently appears in hateful memes that mock familial relationships, highlighting how geographic stereotypes are weaponised for humour and hate.
6 Tag generation module
Figure 2: Schematic of the tag generation module with employed prompts and representative examples explaining each component.
A meme might not always be accompanied by tags, unlike in the case of our dataset. It is therefore important to have an automatic method to assign tags to an arbitrary input meme, which could enhance toxicity detection. Here, we propose a novel generative AI based architecture for obtaining socially relevant tags for a given input meme. Since directly obtaining tags from an input meme is not straightforward, we decompose the task into two steps. First, we generate an intermediate summary describing the input meme; next, we extract keywords from this summary, constituting the final set of tags. This design is inspired by the KeyLLM framework [^12] https://www.maartengrootendorst.com/blog/keyllm/. We next describe each component of the module in detail. Details on employed models and experimental setup are further discussed in Appendix F and H respectively.
Ground truth summary
We curate ground truth summaries that serve as an input for tag generation module discussed later. An overview of the ground truth generation setup is illustrated in Figure 2 (a). We prompt GPT-4o to generate the ground truth summary. We include the following items in the prompt so that a high-quality summary blended with the tags is generated.
(i) The meme’s metadata, e.g., the title and the OCR text.
(ii) The meme’s caption: to generate the image caption, we first apply LaMA image inpainting (Suvorov et al. 2021) to remove any OCR text from the image, and then prompt GPT-4o to produce a caption. This allows for a caption generation that is independent of the text embedded in meme.
(iii) The tags associated with the meme.
(iv) Enhanced context: VLMs being generative models may lack awareness of post-training developments, hence we supplement them with (a) a brief contextual description derived from Google Lens meme search results and (b) an expansion of each of the tags obtained using Google Search API. Details for collecting Google Lens context and its preprocessing are provied in Appendix C.
Manual verification: We randomly sampled 100 memes along with their corresponding GPT-4o generated summaries to conduct a manual verification. Three annotators judged each of the 100 samples where each of them was asked to rate the summaries on a scale of 1 to 10 in terms of the following three criteria: (i) completeness, (ii) fluency, and (iii) grammatical correctness. The evaluation yielded very high average scores of 7.89, 8.48 and 9.15 for completeness, fluency, and grammatical correctness, respectively, hence demonstrating that GPT-4o generated summaries can indeed serve as the ground truth for the pipeline.
Automatic summary generation
We now attempt to automatically generate summaries by fine-tuning various pretrained VLMs – PaliGemma-10B, LLaVA-1.6, and Qwen-2.5. For the fine-tuning, we use the training split noted in Table 2. The fine-tuning prompt queries the VLMs to generate a summary based on the input meme along with the metadata, the image caption and the Google Lens context. Note that we completely hide the tag information from the input. The generated summary is compared with the ground truth summary obtained in the previous section and the generation loss is back propagated. We use PEFT (Mangrulkar et al. 2022) with LoRA adapters (Hu et al. 2021) to support fine-tuning in a resource constrained scenario. The expectation is that this fine-tuning shall enforce the VLMs to learn to include the tags in the generated summary that are already blended in the ground truth summary.
Summary generation performance: We assess the generation performance using the test split noted in Table 2. For each test meme, we compare the generated summary with the corresponding ground truth summary using a suite of standard evaluation metrics, including Bleu score, chrF (character F1), Meteor, Sentence-BERT score, and Rouge. The results are noted in Table 3. We observe that across a majority of the metrics, PaliGemma-10B outperforms the other VLMs. Hence, for the purpose of the final tag extraction module in the next section, we shall use the summaries generated by the PaliGemma-10B model.
Model Bleu chrF MT S-Bert Rouge
PaliGemma-10B 0.1038 40.38 0.3188 77.65 0.2682
Qwen-2.5 0.0821 42.30 0.3156 76.67 0.2443
LLaVA-1.6 0.0292 18.10 0.1257 42.03 0.1117
Table 3: Performance of different VLMs in generating tagged summaries across different metrics. Best results are highlighted. Here, MT: Meteor and S-Bert: Sentence-BERT.
Tag generation
We fine-tune the different VLMs – PaliGemma-10B, LLaVA-1.6, and Qwen-2.5 for generating the final tags. The fine-tuning prompt asks these VLMs to extract a set of tags from the summary of the meme generated by PaliGemma-10B and fed as a part of the prompt. The extracted output tags are compared with the ground truth tags of the meme and the loss is back propagated. The fine-tuning is done again on the training split noted in Table 2. Similarly, we test the performance of the tag generation model on the test split noted in Table 2. In the following we first discuss the baselines and the metrics used for evaluation.
Baselines: We use three state-of-the-art baselines for evaluation. These include (i) Tag2Text (Huang et al. 2024), (ii) RAM (Zhang et al. 2024) and (iii) RAM++ (Huang et al. 2023). Tag2Text, is a vision-language pre-training framework that incorporates image tagging to enhance the learning of visual-linguistic representations. Tag2Text extracts tags directly from associated text, enabling the model to learn an image tagger while simultaneously guiding the vision-language learning process. The RAM model performs image tagging by training on large-scale image-text pairs. An initial model is trained by jointly learning from image captions and tags, supervised respectively by the original textual data and the parsed tags. This is followed by a data refinement engine that generates new tags and removes noisy annotations. The model on this cleaned dataset is retrained and further fine-tuned on a smaller, high-quality set. RAM++ is an open-set image tagging model that effectively harnesses multi-granular textual supervision. RAM++ unifies individual tag-level and global text-level supervision within a single alignment framework. To further enhance tag understanding, it leverages LLMs to transform semantically narrow tag supervision into broader, descriptive tag supervision, thereby enriching the model’s grasp of visual concepts in open-set scenarios. Both RAM and RAM++ models can either use their own trained tags or the open-set vocabulary of tags during inference, resulting in two different variants.
Metrics: We use multiple metrics to evaluate the similarity between the ground truth and the generated tags. These include the semantic similarity [^13] https://sbert.net/docs/sentence˙transformer/usage/semantic˙textual˙similarity.html, BERTScore and ConceptNet similarity [^14] https://conceptnet.io/. For each ground truth tag, we compute the semantic similarity (or BERT score or ConceptNet similarity) with all the generated tags and take the maximum value among these. Next, we average this maximum over all the ground truth tags. We average these averages over all the test data points. In addition, to capture better context, we also expand both the ground truth tag list as well as the generated tag list using the Google Search API and compare the expanded list using the same semantic similarity and BERTScore [^15] ConceptNet is primarily designed for obtaining one-word/short-phrase to one-word/short-phrase relationships. Hence, we have not used it to measure similarity between the expanded tag lists..
Model Tag similarity Exp tag similarity
$S$ BS ConceptNet $S$ BS
PaliGemma-10B 63.9 93.55 51.40 54.5 95.44
Qwen-2.5 59.1 92.65 45.23 48.2 94.77
LLaVA-1.6 56.7 87.89 43.85 45.0 92.88
RAM++ 40.6 90.36 19.43 24.9 92.48
RAM++ Openset 27.8 82.22 8.54 12.8 91.57
RAM 40.5 90.57 19.41 24.8 92.48
RAM Openset 28.3 72.19 7.07 12.2 91.62
Tag2Text 37.3 89.84 17.31 21.2 92.27
Table 4: Performance of the tag generation module. Best results are highlighted. $S$: Semantic similarity, BS: BERTScore, Exp: Expanded.
Image [Uncaptioned image] [Uncaptioned image]
Ground truth trainers, hitler
911 9/11 twin towers impact,
world trade center, jenga,
9/11 truth movement
Our approach hitler, trainers, nazi, just do it 911, twin towers
RAM++ blue, man, shoe , smile building, city
RAM++ Openset
Athletic shoe, Close-up,
Military person, Outdoor shoe
Close-up, Toy block
RAM blue , man , shoe building, city
RAM Openset
Athletic shoe, Black-and-white,
Military person, Outdoor shoe
Toy block
Tag2Text
shoe, picture, uniform,
[…]
building, city,
[…]
Table 5: Comparative examples of tags generated by different models. We provide more examples in Table 7 in the Appendix.
Generation performance: The results of tag generation are presented in Table 4. The top three rows (our scheme) by far outperforms the state-of-the-art baselines in terms all the evaluation metrics. Among our models, PaliGemma-10B has the best tag generation performance. State-of-the-art models fail to capture the meme’s social context and their implicit meaning. They focus primarily on the visible objects in the image (refer Table 5; also refer Table 7 in Appendix for more examples).
7 Toxic meme detection
ToxicTags– In this section we assess the performance of toxic meme detection on the test split for both stages: Stage I and Stage II. Results portray the effectiveness of the tags in few-shot exemplar selection as we demonstrate below (see Table 6). Note that the details of the employed VLMs, embedding generation models, prompt and experimental setup are covered in Appendix FG and H respectively.
Models Shots, Stage $_{\textrm{r}} I$$_{\textrm{im}}$ $_{\textrm{gt}} I$$_{\textrm{pt}}$ $_{\textrm{im}\oplus\textrm{gt}} I$$_{\textrm{im}\oplus\textrm{pt}}$
2, I 51.95 56.94 56.83 58.19 59.66 61.03
2, II 17.59 28.39 24.23 23.79 29.7 29.27
4, I 46.5 62.18 58.51 60.44 62.25 63
IDEFICS-3 4, II 17.58 39.27 35.52 35.45 41.6 41.4
2, I 64.4 66.89 68 67.45 68.56 69.12
2, II 44.29 50.68 53.28 51.41 55.6 54.47
4, I 65.97 70.17 68.86 70.44 70.54 70.34
GPT-4o 4, II 46.10 54.50 52.63 55.27 57.9 57.21
IntMeme I – 45.85 II – 31.6
Table 6: Macro-F1 scores for different few-shot example selection strategies evaluated on the ToxicTags dataset. Further, the final row shows the results for the latest and most competing baseline IntMeme (Hee and Lee 2025) evaluated on the ToxicTags dataset. Best: bold, second best: underline). Overall best result for each stage is highlighted.
Shot selection: In the following we describe the various methods for few-shot selection.
[…]
8 Conclusion
This work makes several key contributions toward advancing research in toxic meme detection and multimodal content moderation. First, we curate a richly annotated dataset of 6,300 real-world meme-based posts through a two-stage labelling process. Second, we enhance contextual understanding by introducing auxiliary metadata including meme titles and, most importantly, social tags. Third, we propose a generative AI-based tag generation module capable of producing socially relevant tags given an arbitrary input meme with no tag information in the first place. Finally, we evaluate the performance of state-of-the-art VLMs under few-shot prompting setups, establishing strong baselines for future research. Collectively, our contributions address key bottlenecks in toxic meme moderation and provide a foundation for building more accurate, context-aware, and socially responsible content moderation systems.
Appendix A Limitations
While this work offers several important contributions, it also has a few limitations. First, our dataset is limited to image-based memes, excluding other emerging modalities such as video and audio, which are increasingly used to disseminate harmful content. Future work will aim to extend this framework to multimodal datasets that better reflect current online communication trends. Second, we do not examine the psychological and emotional impact of hateful memes on viewers. Exposure to such content may contribute to anxiety, depression, or desensitization—an important area that lies beyond the current scope. Addressing this limitation would require collaboration with psychologists, mental health experts, and affected communities. Third, we do not investigate the real-world consequences of online hate memes, particularly their potential to incite offline violence or criminal behavior. Understanding this link between online toxicity and offline harm is a critical direction for future research.
Appendix B Ethics statement
We strictly adhered to the policies of the social media platforms from which the dataset was curated. All memes and associated metadata were manually collected from publicly available content after agreeing to relevant copyright terms, and user anonymity was preserved throughout the process. During annotation, we followed detailed guidelines to safeguard annotators’ mental well-being and conducted two-stage annotation with pilot studies to reduce subjective bias. We acknowledge the potential for bias in both annotation and model predictions; to mitigate this, we employed diverse evaluation metrics and experimental setups to assess model robustness. This study complies with ethical standards for research involving publicly available data, with a focus on transparency, privacy, and minimising harm. Our work aims to support the broader effort to combat online hate while remaining mindful of its own limitations.
Appendix C Platforms
Data curation platform
As stated earlier, we use imgflip.com as the curation platform for our data since it is centred on conversations using memes. We specifically use the streams [^16] https://imgflip.com/streams feature of the platform to collect these memes. The selected streams and their descriptions are outlined as follows:
  • Dark_Humour ($\sim$11k followers) [^17] https://imgflip.com/m/Dark˙humourWelcome to Dark_humour, Imgflip’s premier community for offensive humour. Stream mood: Relax liberals, it’s called dark humour.
  • Memes_Overload ($\sim$9.5k followers) [^18] https://imgflip.com/m/MEMES˙OVERLOADHey there, welcome to MEMES_OVERLOAD, Imgflip’s 4th largest stream for memes! We’re all here to have fun, so make sure to follow the rules, and keep on memeing on.
  • Politics ($\sim$5k followers) [^19] https://imgflip.com/m/politicsHumor and discussion around U.S. and world politics. Criticisms and debates are encouraged, but be constructive and don’t harass anyone.
Figure 3: Post containing meme, title and tags from imgflip platform.
An example cropped post containing the corresponding title and tags is presented in Figure 3.
Google Lens
To provide contextual information from the real world, we provide two different augmentations to our tag generation module: (i) Google Search API based expansion of tags, and (ii) Google Lens based descriptions. While we have discussed in detail about tag expansion in the main content, here we outline the manual collection procedure and preprocessing applied for Google Lens context incorporation.
Collection
Figure 4: Example of manual collection of title and description of most relevant search obtained from Google Lens.
Figure 4 shows an example meme with corresponding similar results obtained from Google Lens. We manually collect this information for each of the 6,300 memes by strictly asking the collectors to keep only the top-most matching result and to adhere to the following format for information storage: Title – Description. We believe that ours is the first of its kind utilisation of Google Lens for improving contextual knowledge.
Preprocessing
For pre-processing, we follow several steps to clean unwanted textual information from the information returned by Google Lens. First, we remove leading and trailing white spaces. Then we remove certain irrelevant URLs appended within the text, which could have generated ambiguity for the tag generation framework and further went on to clean some platform-specific formats such as subreddit references (e.g. r/, u/). Emojis are eliminated using a Unicode-based regex, and numerical metrics (e.g. 2.5K likes, 3M views) & user mentions (e.g. @user) are also eliminated. Finally, attached metadata that do not contribute to meme semantics and non-English scripts (e.g. Chinese, Arabic, Cyrillic, and Devanagari) are also removed. We use ‘Unicode-range matching’ for targeting only English letters, numbers, and symbols while discarding other languages.
Appendix D Details of annotation
Figure 5: A step-by-step flowchart used by annotators to annotate any post.
Figure 5 presents a brief pipeline of the employed annotation process over two stages. The instructions that we designed are given below, and some samples from those provided to annotators are present in Table 8. During the whole annotation process, annotators were asked to adhere to the definitions provided in the subsection D.
  • Each annotator was provided with a separate PDF containing the meme, title, tags and the OCR extracted text from the meme. Along with that, a separate Google sheet was also provided to reduce annotation bias and ensure fairness.
  • In Stage I, they were asked to strictly adhere to the provided definition of toxicity and segregate the memes as toxic or normal. Wherever confusion arose, annotators discussed in our daily scrum and through our Slack workspace.
  • In Stage II, we provided the annotators with toxic memes based on majority voting as per Stage I. Note that before starting Stage II they were fairly aware of the type of content moderation required on these social media platforms. As a first step, they were asked to segregate hateful content, then from the remaining to identify dangerous samples; finally left with offensive samples, which were also verified. All the annotations were performed by strictly adhering to the definition of hateful, dangerous and offensive labels.
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground truth alabama, incest, baby chocolate, rivers, africa
starving cannibal, neon lights,
brothers, cannibalism
Our approach alabama, abortion chocolate, rivers, africa
starving, cannibal,
cannibalism, cannibal
RAM++ anime, brush, girl, person, kiss, love
boy, trench, floor, person, land,
lay, man, mud, puddle, water
barbecue grill, cook, dog, food,
grill, person, man, meat
RAM++ Openset Close-up Jeans Barbecue chicken, Filling station, Ribs
RAM anime, girl, person, kiss
boy, floor, person, land, lay,
man, mud, puddle, water
barbecue, barbecue grill, dog,
food, grill , person, man, meat
RAM Openset ’ ’ Ribs Ribs
Tag2Text
anime, girl, people,
[…]
mud, puddle, water,
[…]
grill, dog, person, man,
[…]
Table 7: Comparative examples of tags generated by different models.
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Title
Does Anyone Else See The
Problem With This
Advertising Sign ?
Cannibal cafe menu If ykyk My friend sent me this
Tags
signs, advertising
hannibal lecter, cannibals,
cannibal, cannibalism
if you know you know, porn
if you know you know,
the owl house, bee movie
OCR extracted
text
imaflip.com ginger
=== S
CANNIBAL CAFE MENU BABY
BACK RIBS FINGER SANDWHICH
OPEN FACE SANDWICH STU STEW
TOE FU SCRAMBLED LEGS
BAKED FRIARS MAC CHEESE
( VEGAN )
Consider John Frazzled
FrazzleMyGimp PIZZA
GUY : You’re total
is \$ 26.34 ME : I can’t
afford that PIZZA GUY :
Well you’$l$ have to pay
some other way ME :
takes out wallet Wait I
forgot I had 30 dollars
PORN DIRECTOR : Cut
According to some ,
this will confuse children
But this is completely
fine HOME TEETH 3
Expert
annotation
hateful dangerous offensive normal
Table 8: Four memes from the expert annotation samples provided to the annotators.
[Uncaptioned image] [Uncaptioned image]
hateful
dangerous
[Uncaptioned image] [Uncaptioned image]
offensive normal
Table 9: Label wise word cloud of tags.
Definitions
hatefulReference(Kiela et al. 2020) – A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease. Attack is defined as violent or dehumanizing (comparing people to non-human things, e.g., animals) speech, statements of inferiority, and calls for exclusion or segregation. Mocking hate crime is also considered hateful.
dangerousReference [^20] https://www.dangerousspeech.org/dangerous-speech – A text, meme or speech which is not hateful but uses any form of expression that can increase the risk of its audience to condone or participate in violence against members of another group will be considered as dangerous.
offensiveReferences(Roy et al. 2023; Mathew et al. 2022) – A text, meme or speech which is neither hateful, nor dangerous but uses abusive slurs or derogatory terms will be considered as offensive.
toxicReference: PerspectiveAPI [^21] https://developers.perspectiveapi.com/s/about-the-api-model-cards?language=en˙US – A rude, disrespectful, or unreasonable comment that is likely to make you leave a discussion.
normal – A meme which is not toxic and follows social norms.
Safety guidelines
Primarily, we performed the following steps to keep our annotators mentally safe with such contents:
(i) Only 50 samples per day were provided to them, wherein we updated the PDF and added corresponding samples in the Google sheet.
(ii) We conducted 15 minutes of daily mental well-being sessions in our daily scrum, by adopting various online activities suggested by WHO [^22] https://www.who.int/news-room/feature-stories/mental-well-being-resources-for-the-public.
(iii) We also asked the annotators to agree to the data source platform’s terms and policies. Further, we strictly ensured that they did not disclose the identity of users in the provided Google sheet.
Appendix E Further analysis of the dataset
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Title Gotta go fast All members get 69% off all items Rouge.exe is getting desperate!
Tags
gotta go fast, sonic
the hedgehog, abortion
human stupidity, ernie and bert,
traffic light
rougeexe, sonic the hedgehog,
free candy van
OCR extracted
text
Just get the back alley abortion and
learn to keep your legs closed ! Welp.
Time to go buy milk . Fast . But
sonic . it’s YOUR baby !
Ernie and Bert want to create
a human trafficking program.
Join today and help make a
difference in your community!
FREE BAT FLAVORED SOUP !
JUST GET IN THE \$#%&ING VAN .
IT ONLY HURTS TIL YOU PASS OUT !
MASKS SAVE LINES
Table 10: Dangerous memes: portrayal of danger through anime/cartoon characters.
tag pairs theme
(‘hitler’, ‘nazi’), (‘hitler’, ‘jews’), (‘hitler’, ’ww2’), (‘hitler’, ‘holocaust’), (‘jews’, ‘nazi’), (‘concentration camp’, ‘nazi’), (‘hitler’, ‘stalin’) Historical events / World War II
(‘9/11’, ‘twin tower’), (‘911’, ‘twin tower’), (‘9/11’, ’muslim’), (‘cow’, ‘muslim’), (‘911 9/11 twin tower impact’, ‘muslim’) 9/11 and Islamophobic references
(‘jesus’, ‘satan’), (‘satan’, ‘the bible’), (‘jesus’, ‘the bible’) Religious contrast / satire
(‘atomic bomb’, ‘hiroshima’), (‘hiroshima’, ‘ww2’), (‘hiroshima’, ‘nuke’) Nuclear warfare / World War II
(‘alabama’, ‘incest’), (‘incest’, ‘sweet home alabama’), (‘alabama’, ‘dad’) Southern US stereotypes, taboo, humour
(‘cannibal’, ‘cannibalism’), (‘fresh’, ‘meat’), (‘cannibalism’, ‘meat’), (‘human’, ‘meat’) Creepy / disturbing themes
(‘mass shooting’, ‘school shooting’), (‘gun’, ‘school shooting’), (‘gun’, ‘school’), (‘gun’, ‘usa’), (‘mass shooting’, ‘usa’), (‘mass shooting’, ‘america’) US gun violence
(‘baby’, ‘dead’), (‘baby’, ‘cannibalism’) Child-related violence / abortion
(‘africa’, ‘water’), (‘africa’, ‘hungry’), (‘africa’, ‘starve’) Humanitarian crises in Africa
(‘black people’, ‘racist’), (‘black’, ‘black life matter’), (‘black’, ‘white’), (‘angry black guy’, ‘lame’) Racism
(‘elmo’, ‘sesame street’), (‘big bird’, ‘sesame street’), (‘ernie’, ‘sesame street’), (‘mayor’, ‘serial killer’), (‘free candy van’, ‘sonic the hedgehog’) Cartoon / anime characters in dark contexts
(‘world war 3’, ‘ww3’), (‘ukraine’, ‘ww3’), (‘gaza’, ‘israel’), (‘monster’, ‘ww3’) Global conflict / war escalation
Table 11: Top co-occurring tag pairs and associated semantic themes.
Related tags: To understand the semantic themes associated with frequently occurring tag pairs, we conduct a co-occurrence analysis. Prior to the analysis, all tags are lemmatised to ensure consistency and improve result accuracy. Our primary objective is to identify how often specific tags co-occur and what thematic contexts they represent within the meme content. The key observations are as follows. Some of the top co-occurring pairs and their themes are listed in Table 11.
Word clusters of tags: Figure 9 presents the word cloud of tags for hateful, dangerous, offensive and normal memes. We can clearly segregate the word clouds for hateful and dangerous, hence signifying our initial observations. For offensive and normal, even though they have similar words, the presence of terms like cursed, nsfw and usage of comic character terms like sesame, street significantly separates them.
Figure 6: Prompts used for Stage I and Stage II classification with few-shot exemplars.
Appendix F Employed models
IDEFICS-3: The latest version of VLM released by HuggingFace team, this is among the few open-source models that are trained with few-shot training objective. We use the instruction fine-tuned version for conducting our experiments – HuggingFaceM4/idefics-9b-instruct.
LLaVA-1.6: An enhanced version of LLaVA (Liu et al. 2023b) and LLaVA-1.5 (Liu et al. 2023a), LLaVA-1.6 has shown significant improvement over its prior models. We use llava-hf/llava-v1.6-mistral-7b-hf checkpoint for running our experiments.
PaliGemma-2: A successor of PaliGemma, PaliGemma-2 presents a series of models of varying size with further improved capabilities due to the incorporation of Gemma-2 LLM instead of Gemma. We use google/paligemma2-10b-pt-224 version from HuggingFace in this work.
Qwen-2.5 VL: Successor of Qwen VL and Qwen-2 VL, this latest iteration introduces streamlined and efficient vision encoder and dynamic resolution for video understanding. We use the instruction-tuned checkpoint from HuggingFace – Qwen/Qwen2.5-VL-7B-Instruct in our experiments.
GPT-4o: One of the first models by OpenAI team to have multimodal capability, GPT-4o (OpenAI 2024) has proved its wide applicability across multiple domains [^23] https://learn.microsoft.com/en-us/azure/ai-services/openai/.
License agreement: We agreed to the terms of usage of all employed VLMs before using them in our work.
Appendix G Employed prompts
Detailed list of prompts for the corresponding variants for Stage-I and Stage-II are provided 6.
Appendix H Experimental setup
(i) Experimental setup for tag generation module:
(a) For generating caption using the LaMa inpainted image, we set max new tokens to 30.
(b) For generating tagged summary we have used maximum number of tokens to 150.
(c) For finetuned summary generator and tag generation module, all the models are finetuned using Trainer API, where we set batch size as 1, gradient accumulation steps set to 1 and a learning rate of 2e-5. We use AdamW Torch optimizer with a $\beta_{2}$ value of 0.999 and a weight decay of 1e-6. Transformer modules including q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj and down_proj, with a rank value of 8 and task type set to causal language modeling, are finetuned using LoRA-based parameter-efficient fine-tuning. We use 4-bit quantization with bfloat16 compute precision through the BitsAndBytesConfig for efficient memory usage and faster training. To stabilize the initial learning phase, we set a warmup step of two and the number of training epochs to three. To control the maximum sequence length, we set max length to 512, while truncation and padding are also applied.
(ii) Few-shot prompting: To control the randomness of the generated output, we set the temperature to 0.001 for all models used in Stage I and II. The max new tokens was set to 30 and all other parameters were kept as default.

### v2

@@
-ToxicTags: Decoding Toxic Memes with Rich Tag Annotations
+StemTox: From Social Tags to Fine-Grained Toxic Meme Detection via Entropy-Guided Multi-Task Learning
 Abstract
-The 2025 Global Risks Report [^1] https://reports.weforum.org/docs/WEF˙Global˙Risks˙Report˙2025.pdf identifies state-based armed conflict and societal polarisation among the most pressing global threats, with social media playing a central role in amplifying toxic discourse.
 Memes, as a widely used mode of online communication, often serve as vehicles for spreading harmful content.
-However, limitations in data accessibility and the high cost of dataset curation hinder the development of robust meme moderation systems.
-To address this challenge, in this work, we introduce a first-of-its-kind dataset of 6,300 real-world meme-based posts annotated in two stages: (i) binary classification into toxic and normal, and (ii) fine-grained labelling of toxic memes as hateful, dangerous, or offensive.
+However, limitations in data accessibility and the high costs of dataset curation hinder the development of robust meme moderation systems.
+To address this challenge, in this work, we introduce a first-of-its-kind dataset – ToxicTags consisting of 6,300 real-world meme-based posts annotated in two stages: (i) binary classification into toxic and normal, and (ii) fine-grained labelling of toxic memes as hateful, dangerous, or offensive.
 A key feature of this dataset is that it is enriched with auxiliary metadata of socially relevant tags, enhancing the context of each meme.
-In addition, we propose a tag generation module that produces socially grounded tags, because most in-the-wild memes often do not come with tags.
+In addition, we propose a novel entropy guided multi-tasking framework – StemTox – that integrates the generation of socially grounded tags with a robust classification framework.
 Experimental results show that incorporating these tags substantially enhances the performance of state-of-the-art VLMs in toxicity detection tasks.
@@
 In response to these challenges, vision language models (VLMs) have recently gained traction as powerful tools for content moderation.
-These models are capable of jointly analyzing visual and textual elements to grasp the nuanced context of memes and can provide detailed justifications for their classifications (Qu, Backes, and Zhang 2025).
-Despite these advances, recent research highlights the limitations of VLMs in accurately detecting hateful memes (Rizwan et al.
-2025), underscoring the urgent need to bridge these performance gaps.
+These models are capable of jointly analyzing visual and textual elements to grasp the nuanced context of memes and can provide detailed justifications for their classifications Qu et al.
+(2025).
+Despite these advances, recent studies have highlighted the limitations of VLMs in accurately detecting hateful memes Rizwan et al.
+(2025), underscoring the urgent need to bridge these performance gaps.
 Similarly, a recent blog post [^2] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes/ by Meta highlighted the challenges and complexities inherent in their current content moderation frameworks.
-Dataset scarcity: Datasets serve as the foundational fuel for generative AI models.
+Datasets serve as the foundational fuel for generative AI models.
 However, researchers increasingly face significant obstacles in curating such datasets, primarily due to the high costs of manual annotation and the limitations imposed on large-scale crawling of social media platforms.
-These challenges have resulted in datasets that are either manually constructed (Kiela et al.
-2020; Lin et al.
-2024) – often failing to fully capture the richness of human creativity – or narrowly scoped to specific targets or events (Pramanick et al.
-2021a, b; Fersini et al.
-2022).
-Moreover, this scarcity of comprehensive datasets has hindered efforts to simulate the complexity of social media ecosystems across a broad taxonomy of content labels.
-This is in contrast to textual hate speech research, where more extensive taxonomical explorations exist (Sachdeva et al.
-2022).
-Our contributions: To address the aforementioned limitations, we make the following contributions.
-We introduce a novel, highly diverse, real-world meme dataset enriched with human-annotated contextual tags – a critical yet often overlooked feature in social media content.
+These challenges have resulted in datasets that are either manually constructed Kiela et al.
+(2020); Lin et al.
+(2024) – often failing to fully capture the richness of human creativity – or narrowly scoped to specific targets or events Pramanick et al.
+(2021a, b); Fersini et al.
+(2022).
+Moreover, this scarcity of comprehensive datasets has hindered efforts to build robust toxicity detection systems for multimodal social media content.
+This is in contrast to textual hate speech research, where more extensive taxonomical explorations exist Sachdeva et al.
+(2022).
+In this work, we present, for the first time, a diverse real-world dataset of memes labelled across four nuanced categories – toxic, hateful, dangerous, offensive – alongside normal, moving beyond the traditional binary classification paradigm predominantly used in prior studies (refer to Table 1).
+In addition, each meme in the dataset is also labelled with a set of social tags that contextually ground the meme.
+Finally, we introduce StemTox, a novel, socially aware entropy-guided multi-tasking framework, opening new avenues for open-set image tagging and robust toxicity classification.
+Our contributions are noted below.
+The key contributions of our work are as follows: ☞ We introduce a novel, highly diverse, real-world meme dataset ToxicTags enriched with human-annotated contextual tags – a critical yet often overlooked feature in social media content.
 Unlike prior datasets, it consists of only real-world memes with no restrictions based on specific targets or events (see Section 3).
 The final dataset comprises 6,300 annotated memes, capturing a wide spectrum of online discourse.
-We design a rigorous two-stage human annotation pipeline.
-In the first stage, memes are classified as either toxic or normal.
+☞ We design a rigorous two-stage human annotation pipeline.
+In the first stage, memes are annotated as either toxic or normal.
 In the second stage, toxic memes are further categorized into one of three fine-grained classes: hateful, dangerous, or offensive.
 These categories have been distilled through an iterative annotation process, revealing that a four-class taxonomy is appropriately expressive for moderating a wide range of social media content.
-This taxonomy can help reduce misclassifications in existing moderation pipelines.
-We propose a novel tag generation framework (see Section 6) significantly departing from existing approaches.
-Leveraging real-world context, our system integrates Google Search API and Google Lens to produce high-quality, socially grounded tags for memes.
-We perform few-shot prompting of VLMs using various schemes to select the samples.
-Our experiments demonstrate that the inclusion of contextual tags significantly enhances model performance in meme classification tasks.
-Finally, we report the performance of our method on a standard hate meme detection dataset.
-Note that this dataset does not have the tag information in the first place; therefore, we use our tag generation module to predict relevant tags for a query image.
-Inclusion of these tags in the scheme for sampling the few-shot examples results in a better performance compared to the case where the tags are absent.
+This taxonomy can help reduce misclassifications in existing toxicity detection pipelines.
+☞ We introduce StemTox, a novel framework that significantly advances the state-of-the-art by generating socially grounded, high-quality tags and improving meme classification performance.
+Our experiments demonstrate that incorporating contextual tags into the multitasking system leads to significant performance gains in meme classification tasks.
+☞ Finally, we report the performance of our method on two standard benchmark hate meme detection datasets and compare it against five baseline approaches.
 2 Related Works
-Hate meme detection: Rapid surge of hate around the globe (Arcila Calderón et al.
-2024) in the recent past years, with memes acting as a major source of fuel, has led to the curation of multiple hateful memes dataset (Kiela et al.
-2020; Lin et al.
-2025).
-Similarly, there has been extensive research in the field to build robust content moderation frameworks (Rizwan et al.
-2025; Das and Mukherjee 2023b; Prakash et al.
-2023; Cao, Lee, and Jiang 2024; Cao et al.
-2023), with some works on low-resource languages (Das and Mukherjee 2023a; Kumari et al.
-2024).
-Despite such rapid developments, current datasets are either limited to manually curated memes or focus only on a subset of events (Pramanick et al.
-2021a, b; Chen et al.
-2023).
-dataset
-event/target
-independent?
-real-world
-memes?
-post’s contextual
-information?
-two stage
-annotation?
-size labels
-FHM (Kiela et al. 2020) $\sim$10,000 hateful, not-hateful
-MAMI (Fersini et al. 2022) $\sim$10,000 misogynistic, not-misogynistic
-HARM-C (Pramanick et al. 2021a) $\sim$3,544 very harmful, partially harmful, harmless
-HARM-P (Pramanick et al. 2021b) $\sim$3,552 very harmful, partially harmful, harmless
-UA-RU Conflict  (Thapa et al. 2022) $\sim$5,680 hateful, not-hateful
-CrisisHateMM (Bhandari et al. 2023) $\sim$4,723 hateful, not-hateful
-RUHate-MM (Thapa et al. 2024) $\sim$20,675 hateful, not-hateful
-ToxicTags (ours) $\sim$6,300
-I– toxic, normal
-II– hateful, dangerous, offensive, normal
-Table 1: Comparison of our dataset with existing datasets on toxic memes. Here post’s contextual information refers to the title or tags associated with the meme (as in ToxicTags).
-Tags: Prior research has extensively explored automated tag generation for images (Huang et al.
-2024; Zhang et al.
-2024; Dai et al.
-2023), leading to the development of several dedicated models and datasets.
+Hate meme detection: Rapid surge of hate around the globe Arcila Calderón et al.
+(2024) in the recent past years, with memes acting as a major source of fuel, has led to the curation of multiple hateful memes dataset Kiela et al.
+(2020); Lin et al.
+(2025).
+Similarly, there has been extensive research in the field to build robust content moderation frameworks Rizwan et al.
+(2025); Das and Mukherjee (2023b); Prakash et al.
+(2023); Cao et al.
+(2024a, 2023b), with some works on low-resource languages Das and Mukherjee (2023a); Kumari et al.
+(2024).
+Despite such rapid developments, current datasets are either limited to manually curated memes or focus only on a subset of events Pramanick et al.
+(2021a, b); Chen et al.
+(2023).
+dataset event / target independent? real world memes? post’s contextual information? two stage annotation? size labels
+FHM Kiela et al. (2020) $\sim$10k hateful, not-hateful
+MAMI Fersini et al. (2022) $\sim$10k misogynistic, not-misogynistic
+HARM-C Pramanick et al. (2021a) $\sim$3,544 very/partially harmful, harmless
+HARM-P Pramanick et al. (2021b) $\sim$3,552 very/partially harmful, harmless
+UA–RU Conflict Thapa et al. (2022) $\sim$5,680 hateful, not-hateful
+CrisisHateMM Bhandari et al. (2023) $\sim$4,723 hateful, not-hateful
+RUHate-MM Thapa et al. (2024) $\sim$20,675 hateful, not-hateful
+ToxicTags $\sim$6,300 stage I: toxic, normal stage II: hateful, dangerous, offensive, normal
+Table 1: Comparison of ToxicTags with existing datasets for toxic meme detection. Post-contextual information includes metadata such as titles or tags associated with the meme.
+Tags: Prior research has extensively explored automated tag generation for images Huang et al.
+(2024); Zhang et al.
+(2024); Dai et al.
+(2023), leading to the development of several dedicated models and datasets.
 However, to the best of our knowledge, no existing work addresses the specific challenge of tag generation for memes – visual artefacts that often combine text and imagery in culturally nuanced, context-dependent ways.
-To fill this gap, we curate a dedicated meme-tagging dataset and introduce a novel framework built around KeyLLM [^4] https://maartengr.github.io/KeyBERT/guides/keyllm.html to generate socially relevant and context-aware tags from memes (see Section 6).
-This work: We present, for the first time, a diverse real-world dataset of memes labelled across four nuanced categories – toxic, hateful, dangerous, offensive – alongside normal, moving beyond the traditional binary classification used in prior studies (refer to Table 1).
-In addition, we introduce a novel, socially aware tag generation module, opening new avenues in open-set image tagging.
-Finally, we employ these tags to improve the performance of toxic meme detection.
-3 Dataset
+Present work: We present a real-world meme dataset with fine-grained toxicity labels beyond binary classification (Table 1).
+In addition, each meme is annotated with a set of socially aware tags.
+We also introduce a StemTox for open-set image tagging and robust classification.
+3 The ToxicTags dataset
 Collection: We source real-world social media memes from [https://imgflip.com], specifically utilizing its streams section [^3] https://imgflip.com/streams to crawl meme-centric posts.
 The platform was chosen for two primary reasons: (i) its strong emphasis on social media-style interactions where memes serve as a central mode of communication, and (ii) its wide variety of user-generated content, allowing us to collect memes without any target-specific or event-specific filtering, thereby closely mirroring real-world social media environments (refer to Table 1 for comparison with existing datasets).
-For our dataset, we select three high-volume and thematically rich streams [^6] https://imgflip.com/m/Dark˙humour, https://imgflip.com/m/MEMES˙OVERLOAD, https://imgflip.com/m/politicsDark_Humour, Memes_Overload, and Politics based on their popularity and relevance to diverse meme content.
-This selection has been made through manual inspection of stream descriptions and content samples (see Appendix C for details).
+For our dataset, we select three high-volume and thematically rich streams—Dark_Humour [^4] https://imgflip.com/m/Dark_humour, Memes_Overload [^5] https://imgflip.com/m/MEMES_OVERLOAD, and Politics [^6] https://imgflip.com/m/politics—based on their popularity and relevance to diverse meme content.
+We selected these streams through manual inspection of their descriptions and content samples(see Appendix C for details).
 Further, to ensure that the memes included are socially engaging and contextually rich, we manually review a subset of posts and observe that memes with fewer than two comments are often unengaging or irrelevant to a broader audience.
 Hence, we retain only those memes that received at least two comments.
-In total, we manually curate 37,072 memes across the selected streams over a period of approximately one month.
+In total, we manually curated 37,072 memes across the selected streams over a period of approximately one month.
 The memes were downloaded using a browser-based image downloader extension [^7] https://chromewebstore.google.com/detail/download-all-images/ifipmflagepipjokmbdecpmjbibjnakm?hl=en.
@@
 To ensure robustness of OCR extracted text, we perform manual verification of 100 randomly chosen samples and find nearly 98 to be correctly identified; thus depicting the robustness of the tool.
-Further, we also store the bounding box of OCR and use it later for generating tags (refer to Section 6).
 Pre-processing: Before providing the meme along with its metadata for annotation, we perform the following pre-processing steps.
@@
 Subsequently, we obtain a rich set of 7,209 unique and socially relevant tags from an initial list of 9,664 unique tags.
-Agreement: We strictly adhere to the privacy and copyright regulations of the platform [^10] https://imgflip.com/terms and therefore collected our data only from the publicly available posts.
+Agreement: We strictly adhered to the privacy and copyright regulations of the platform [^10] https://imgflip.com/terms and therefore collected our data only from the publicly available posts.
 To maintain the privacy of users, we did not store any information that could potentially compromise their anonymity.
@@
 Two staged annotation: We adopt a two stage annotation process – (i) classification of each meme into toxic or normal, and (ii) further bifurcation of toxic memes into offensive, dangerous or hateful.
-The two-stage process is specifically adopted to mitigate annotation errors as suggested in (Rizwan et al.
-2025).
-Further, we also conduct a pilot study at each stage before concluding the final annotations.
-For both stages, we perform three annotations per sample by three different annotators to minimise annotation bias, given that this is a highly subjective task (Mathew et al.
-2022).
+The two-stage process is specifically adopted to mitigate annotation errors as suggested in Rizwan et al.
+(2025).
+Figure 1: A step-by-step flowchart used by annotators to annotate any post.
+Figure 1 presents a brief pipeline of the employed annotation process over two stages.
+The instructions that we designed are given below, and some samples from those provided to annotators are present in Table 10.
+During the whole annotation process, annotators were asked to strictly adhere to the label definitions (refer Appendix D).
+Further, we also conducted a pilot study at each stage before progressing to the final annotations.
+For both stages, we perform three annotations per sample by three different annotators to minimise annotation bias, given that this is a highly subjective task Mathew et al.
+(2022).
 We use the standard definition of labels that are effectively used for practical applications.
-In the upcoming subsections, we discuss each stage separately and further present detailed annotation instructions with corresponding definitions of the labels in Appendix D (also see Figure 5 in the Appendix).
-Annotator details: We initially selected 25 annotators through an open call in a university setting.
-The minimum criteria that we set are (i) the enrollment of the candidate in the second year of their bachelor’s program, and (ii) familiarity of the candidate with diverse social media content.
-Among these, five are experienced researchers having exposure to this field, and the rest 20 candidates are engineering students in their sophomore years.
-All the annotators are within the age range of 21-27 years, and the whole annotation process has been conducted under the close supervision of two senior researchers.
-To help the annotators with their regular queries and to make sure their mental health is not affected, we conducted daily scrum and further made a Slack workspace for easier collaboration.
-In particular, we asked all the participating annotators to sign an agreement before starting the annotation, wherein we specifically focus on their mental well-being, given the nature of the work.
-We detail out the safety guidelines in Appendix D; further, as an extra safety measure, we rolled out a maximum of only 50 samples per day to each annotator.
-Annotation tool: Separate PDFs for memes with all required metadata (as shown in Figure 5 in the Appendix) and a separate Google sheet have been provided to each annotator to reduce bias and ensure fair annotation.
-Further in our scrum with the annotators, we assisted them by discussing their queries to provide them further insights so as to maintain a high quality of annotation.
+In the upcoming subsections, we discuss each stage separately and further present detailed annotation instructions with corresponding definitions of the labels in Appendix D.
+Annotator details: To ensure a high-quality annotation process, we shortlisted 25 annotators based on the following minimum criteria: (i) candidates having professional experience and prior work aligned with our research, and (ii) familiarity with diverse social media content.
+We additionally required annotators to pass a qualification task curated by the supervising researchers to assess label comprehension and consistency.
+The entire annotation process was conducted under the close supervision of two PhD researchers.
+We detail the safety guidelines in Appendix D; further, as an extra safety measure, we rolled out a maximum of only 50 samples per day to each annotator.
 Annotation codebook: We also provide detailed annotation instructions, accompanied by 25 samples that have been manually selected and annotated by the two supervising researchers.
@@
 For the toxic label, we strictly follow the definition used by the Perspective API, given its widespread adoption in production settings.
-For the hateful label, we adhere to the Facebook hate meme definition (Kiela et al.
-2020), which has become the de facto standard.
+For the hateful label, we adhere to the Facebook hate meme definition Kiela et al.
+(2020), which has become the de facto standard.
 For the dangerous label, we follow the annotation guidelines available at the following source  [^11] https://www.dangerousspeech.org/dangerous-speech.
-Stage I: Binary labelling of the dataset
+4.1 Stage I: Binary labelling of the dataset
 Annotation: The full set of 6,300 samples has been evenly distributed among the 12 selected annotators (out of the 25 recruited).
-Each post has been independently annotated by three annotators.
-We achieve a final inter-annotator agreement score of 0.8176, as measured by Fleiss’ $\kappa$.
-Notably, for a subjective task (Mathew et al.
-2022) such as ours, this level of agreement is considered relatively high, thereby validating the quality of our annotations.
+As stated earlier, we employ three annotators per sample which enhances annotation diversity.
+We achieve a final inter-annotator agreement score of 0.753, as measured by Fleiss’ $\kappa$.
+Notably, for a subjective task Mathew et al.
+(2022) such as ours, this level of agreement is considered relatively high, thereby validating the quality of our annotations.
 Each annotator has been paid USD 40 for this task which is much above the minimum wage in the annotators’ country.
 Label assignment: The final label for each sample is determined based on the majority vote of the three annotations, i.e., the label with at least two agreements is considered to be the final label.
-In this stage, we finally obtained 1,392 normal and 4908 toxic samples (refer to Table 2 for complete statistics).
-Stage II: Fine-grained labelling
+In this stage, we finally obtained 1,446 normal and 4,854 toxic samples (refer to Table 2 for complete statistics).
+Figure 2: Frequently occurring tags in the toxic, hateful, offensive, and normal classes.
+4.2 Stage II: Fine-grained labelling
 Annotation: As a pilot step, we randomly sampled 250 toxic memes from the previous stage and had the expert annotators categorise them into three distinct labels.
@@
 Upon verification, we found that the annotators were performing satisfactorily, and thus proceeded with the final annotation with this group.
-Subsequently, the set of 4,908 toxic samples was categorised into three classes: hateful, dangerous, and offensive.
+Subsequently, the set of 4854 toxic samples was categorised into three classes: hateful, dangerous, and offensive.
 As before, each sample received three independent annotations.
-The annotators achieved a Fleiss’ $\kappa$ agreement score of 0.8197 – a slight improvement over Stage I – highlighting the value of multi-stage and multi-pilot studies in enhancing annotation quality.
+The annotators achieved a Fleiss’ $\kappa$ agreement score of 0.793 – a slight improvement over Stage I – highlighting the value of multi-stage and multi-pilot studies in enhancing annotation quality.
 Label assignment: The final label is determined based on the majority vote of the annotators for each sample.
-When no label receives more than one vote (which means equal vote to hateful, dangerous and offensive for that particular sample), we exclude that sample from the dataset and put it in undecided category.
+When no label receives more than one vote (which means equal vote to hateful, dangerous and offensive for that particular sample).
+For undecidable cases, final annotation was carried out by our expert researchers.
 The overall dataset statistics are noted in Table 2.
@@
 stage label train test total
-I & II normal 1243 149 1392
-I toxic 4057 851 4908
-hateful 1475 343 1818
-dangerous 1788 375 2163
-offensive 653 122 775
-II undecided 141 11 152
+I & II normal 1297 149 1446
+I toxic 4003 851 4854
+II hateful 1470 282 1752
+dangerous 1847 472 2319
+offensive 686 97 783
 total 5300 1000 6300
-Table 2: Data statistics showing the distribution of the data points across the classes.
+Table 2: Dataset statistics showing the distribution of samples across classes.
+tag pairs theme
+(‘hitler’, ‘nazi’), (‘hitler’, ‘jews’), (‘hitler’, ’ww2’), (‘hitler’, ‘holocaust’), (‘jews’, ‘nazi’), (‘concentration camp’, ‘nazi’), (‘hitler’, ‘stalin’) Historical events / World War II
+(‘9/11’, ‘twin tower’), (‘911’, ‘twin tower’), (‘9/11’, ’muslim’), (‘cow’, ‘muslim’), (‘911 9/11 twin tower impact’, ‘muslim’) 9/11 and Islamophobic references
+(‘jesus’, ‘satan’), (‘satan’, ‘the bible’), (‘jesus’, ‘the bible’) Religious contrast / satire
+(‘atomic bomb’, ‘hiroshima’), (‘hiroshima’, ‘ww2’), (‘hiroshima’, ‘nuke’) Nuclear warfare / World War II
+(‘alabama’, ‘incest’), (‘incest’, ‘sweet home alabama’), (‘alabama’, ‘dad’) Southern US stereotypes, taboo, humour
+(‘cannibal’, ‘cannibalism’), (‘fresh’, ‘meat’), (‘cannibalism’, ‘meat’), (‘human’, ‘meat’) Creepy / disturbing themes
+(‘mass shooting’, ‘school shooting’), (‘gun’, ‘school shooting’), (‘gun’, ‘school’), (‘gun’, ‘usa’), (‘mass shooting’, ‘usa’), (‘mass shooting’, ‘america’) US gun violence
+(‘baby’, ‘dead’), (‘baby’, ‘cannibalism’) Child-related violence / abortion
+(‘africa’, ‘water’), (‘africa’, ‘hungry’), (‘africa’, ‘starve’) Humanitarian crises in Africa
+(‘black people’, ‘racist’), (‘black’, ‘black life matter’), (‘black’, ‘white’), (‘angry black guy’, ‘lame’) Racism
+(‘elmo’, ‘sesame street’), (‘big bird’, ‘sesame street’), (‘ernie’, ‘sesame street’), (‘mayor’, ‘serial killer’), (‘free candy van’, ‘sonic the hedgehog’) Cartoon / anime characters in dark contexts
+(‘world war 3’, ‘ww3’), (‘ukraine’, ‘ww3’), (‘gaza’, ‘israel’), (‘monster’, ‘ww3’) Global conflict / war escalation
+Table 3: Top co-occurring tag pairs and associated semantic themes.
+Figure 3: Proposed StemTox framework.
 5 Analysis of the dataset
-Figure 1: Frequently occurring tags in the toxic, hateful, offensive, and normal classes.
-One of the most unique features of our dataset is the tags associated with each meme.
-This allows us to perform certain interesting analysis that we report below.
-Some of the most frequent tags we find in the train and the test splits are illustrated in Figure 1.
+One of the most unique features of our dataset is the tags associated with each meme. This allows us to perform certain interesting analysis that we report below.
+Frequent tags: Some of the most frequent tags we find in the train and the test splits are illustrated in Figure 2.
 We report the top 30 tags from the toxicity class as well as the top 10 tags each from the hateful, dangerous, offensive and normal classes, respectively.
-Some properties of the top tags are enlisted below (see Table 11 in the Appendix for more details).
+The key observations are as follows.
 (i) The distribution of the top tags across the train and the test splits is very similar, indicating that our strategy of splitting the data is effective.
@@
 We also observe that for the dangerous class, some tags that appear benign at the surface level and are primarily associated with children’s media show up.
-These include ‘ernie and bert’, ‘sesame street’, and ‘sonic the hedgehog’ which are used to render a comic angle to the dangerous posts (see Table 10 in the Appendix for more examples).”
+These include ‘ernie and bert’, ‘sesame street’, and ‘sonic the hedgehog’ which are used to render a comic angle to the dangerous posts (see Table 12 in the Appendix for more examples)."
 (iii) We observe that users are generally reluctant to reshare content involving serious violence, suicide, or murder, as indicated by the relatively low frequency of the ‘repost’ tag in the dangerous category compared to its prevalence in other categories.
@@
 (vi) One notable observation concerns the use of the tag ‘alabama’, which frequently appears in hateful memes that mock familial relationships, highlighting how geographic stereotypes are weaponised for humour and hate.
-6 Tag generation module
-Figure 2: Schematic of the tag generation module with employed prompts and representative examples explaining each component.
+Related tags: To understand the semantic themes associated with frequently occurring tag pairs, we conduct a co-occurrence analysis.
+Prior to the analysis, all tags are lemmatised to ensure consistency and improve result accuracy.
+Our primary objective is to identify how often specific tags co-occur and what thematic contexts they represent within the meme content.
+Some of the top co-occurring pairs and their themes are listed in Table 3.
+Our study reveals that a large portion of frequent tag pairs relate to the Historical events/World War II theme, which indicates that memes are often coded around historical events for satire or shock value.
+From Table 3, we also observe that, to create humour or evoke strong reactions memes often engage with religious contrast/satire themes and controversial cultural commentary, such as 9/11 and Islamophobic references.
+Several pairs relate to themes such as nuclear warfare/World War II, child-related violence or abortion, creepy and disturbing content, and U.S.
+gun violence, indicating that memes frequently present serious or sensitive issues through humour or provocation.
+Themes such as Southern US stereotypes, taboo humour, humanitarian crises in Africa, and racism encode racial, regional, or humanitarian stereotypes, reflecting controversial social issues.
+Cartoon/anime character themes are often used in dark contexts.
+6 The StemTox framework
+6.1 Entropy-guided layer selection objective
+Past approaches in toxic meme detection have focused primarily on single-task setups and rely either on zero-/few- shot prompting, or perform low-resource fine-tuning using stratetgies like PEFT/LoRA.
+Recently, multimodal research has evolved, and some works like He et al.
+(2025); Luo et al.
+(2024) propose a deeper understanding of the model’s internal representations and semantic alignment.
+A recent empirical evidence presented in Chen et al.
+(2025) suggests that deep representational alignment between vision and language modalities peaks at intermediate depths before undergoing representational drift in terminal layers.
+Taking cues from these works, we propose a novel token confidence based approach for simultaneous tag generation and toxic meme classification.
+Below, we detail our approach.
+Figure 3 illustrates the complete workflow of StemTox.
+We consider a multimodal model $\mathcal{M}$ that takes a text prompt $\mathcal{T}$ and an image $\mathcal{I}$ as input and consists of $L$ stacked layers.
+Given an input sequence of fixed length, the model produces hidden representations across layers.
+Let $l\in\{1,\dots,L\}$ denote the layer index, $t\in\{1,\dots,T\}$ denote the token position in the input sequence, and $v\in V$ denote a token from the model vocabulary $V$.
+The hidden state $\mathbf{h}_{l,t}\in\mathbb{R}^{d_{h}}$ corresponds to token $t$ at layer $l$, as depicted in the figure, where $d_{h}$ represents the hidden dimension of the model.
+To process the model’s internal state, we extract token embeddings from all layers and project them into the model’s vocabulary space using the pretrained language model head $\mathbf{W}_{\text{vocab}}\in\mathbb{R}^{d_{h}\times|V|}$.
+This projection produces $\mathbf{w}_{l,t}$, which indicates the representation of the $t$-th token at layer $l$ in the vocabulary V:
+$\mathbf{w}_{l,t}=\mathbf{h}_{l,t}\,\mathbf{W}_{\text{vocab}}.$ (1)
+A softmax function is applied to $\mathbf{w}_{l,t}$ to obtain the confidence distribution $p_{l,t}$ over the vocabulary for the $t$-th token at layer $l$:
+$p_{l,t}=\mathrm{softmax}\!\left(\mathbf{w}_{l,t}\right),$ (2)
+To measure the confidence of the representation at layer $l$ for token $t$, we compute the Shannon entropy Vajapeyam (2014) $e_{l,t}$, which represents the entropy of token $t$ at layer $l$, as follows:
+$e_{l,t}=-\sum_{v\in V}p_{l,t}(v)\,\log p_{l,t}(v).$ (3)
+This entropy measure enables the dynamic selection of the least random token embedding across layers by selecting the representation corresponding to the minimum entropy, enabling the classification head to extract features from the most certain internal representations $\mathbf{h}_{\text{eff}}$:
+$\mathbf{h}_{\text{eff}}=\mathrm{Aggregate}\!\left(\left\{\mathbf{h}_{l^{*},t}\;\middle|\;l^{*}=\arg\min_{l\in\mathcal{L}}e_{l,t}\right\}_{t=1}^{T}\right).$ (4)
+Finally, the confident sequence representation $\mathbf{h}_{\text{eff}}$ is passed through the classification head to predict the class logits.
+The classification head consists of a sequence compression layer that compresses the sequence and projects it to a fixed-dimensional representation, which is then passed through an Multi-Layer Perceptron(MLP) for final classification:
+$\textrm{classification label}=\mathrm{MLP}\!\left(\mathrm{Compress}\!\left(\mathbf{h}_{\text{eff}}\right)\right)$ (5)
+We employ a weighted cross-entropy loss for the classification task, which is depicted as $\mathcal{L}_{\text{cls}}$ in the figure.
+To address class imbalance, class weights are computed following the weighting scheme described in King and Zeng (2001) [^12] https://scikit-learn.org/stable/modules/generated/sklearn.utils.class_weight.compute_class_weight.html.
+approach model task/ shot toxicity detection tag similarity
+stage I stage II chrF met $S$
+mF1 acc mF1 acc
+FS (TS) GPT 2-shot 61.26 75.37 48.59 51.70
+4-shot 63.37 78.87 51.78 55.76
+LHL PaliGemma uni 61.53 70.40 45.92 57.80
+multi 70.48 82.20 61.54 69.20 0.4782 0.26 0.6220
+LLaVA uni 61.58 74.60 44.69 58.70
+multi 65.03 76.00 54.56 64.40 0.2585 0.1050 0.4540
+StemTox PaliGemma uni 60.85 74.90 42.78 52.90
+multi 72.55 83.90 64.66 72.40 0.4872 0.265 0.626
+LLaVA uni 62.19 77.20 51.19 61.60
+multi 67.24 77.70 57.99 66.90 0.2648 0.1070 0.4470
+Table 4: Comparative results (reported out of 100) for toxicity detection and tag similarity across different approaches.
+The best results are highlighted.
+FS (TS): few-shot based on tag similarity; LHL: last hidden layer; mF1: Macro-F1; acc: Accuracy; chrF: chr-F; met: Meteor; $S$: Semantic-Similarity.
+(‘–’ indicates the task is not supported by methods.)
+6.2 Multitask learning objective
 A meme might not always be accompanied by tags, unlike in the case of our dataset.
 It is therefore important to have an automatic method to assign tags to an arbitrary input meme, which could enhance toxicity detection.
-Here, we propose a novel generative AI based architecture for obtaining socially relevant tags for a given input meme.
-Since directly obtaining tags from an input meme is not straightforward, we decompose the task into two steps.
-First, we generate an intermediate summary describing the input meme; next, we extract keywords from this summary, constituting the final set of tags.
-This design is inspired by the KeyLLM framework [^12] https://www.maartengrootendorst.com/blog/keyllm/.
-We next describe each component of the module in detail.
-Details on employed models and experimental setup are further discussed in Appendix F and H respectively.
-Ground truth summary
-We curate ground truth summaries that serve as an input for tag generation module discussed later.
-An overview of the ground truth generation setup is illustrated in Figure 2 (a).
-We prompt GPT-4o to generate the ground truth summary.
-We include the following items in the prompt so that a high-quality summary blended with the tags is generated.
-(i) The meme’s metadata, e.g., the title and the OCR text.
-(ii) The meme’s caption: to generate the image caption, we first apply LaMA image inpainting (Suvorov et al.
-2021) to remove any OCR text from the image, and then prompt GPT-4o to produce a caption.
-This allows for a caption generation that is independent of the text embedded in meme.
-(iii) The tags associated with the meme.
-(iv) Enhanced context: VLMs being generative models may lack awareness of post-training developments, hence we supplement them with (a) a brief contextual description derived from Google Lens meme search results and (b) an expansion of each of the tags obtained using Google Search API.
-Details for collecting Google Lens context and its preprocessing are provied in Appendix C.
-Manual verification: We randomly sampled 100 memes along with their corresponding GPT-4o generated summaries to conduct a manual verification.
-Three annotators judged each of the 100 samples where each of them was asked to rate the summaries on a scale of 1 to 10 in terms of the following three criteria: (i) completeness, (ii) fluency, and (iii) grammatical correctness.
-The evaluation yielded very high average scores of 7.89, 8.48 and 9.15 for completeness, fluency, and grammatical correctness, respectively, hence demonstrating that GPT-4o generated summaries can indeed serve as the ground truth for the pipeline.
-Automatic summary generation
-We now attempt to automatically generate summaries by fine-tuning various pretrained VLMs – PaliGemma-10B, LLaVA-1.6, and Qwen-2.5.
-For the fine-tuning, we use the training split noted in Table 2.
-The fine-tuning prompt queries the VLMs to generate a summary based on the input meme along with the metadata, the image caption and the Google Lens context.
-Note that we completely hide the tag information from the input.
-The generated summary is compared with the ground truth summary obtained in the previous section and the generation loss is back propagated.
-We use PEFT (Mangrulkar et al.
-2022) with LoRA adapters (Hu et al.
-2021) to support fine-tuning in a resource constrained scenario.
-The expectation is that this fine-tuning shall enforce the VLMs to learn to include the tags in the generated summary that are already blended in the ground truth summary.
-Summary generation performance: We assess the generation performance using the test split noted in Table 2.
-For each test meme, we compare the generated summary with the corresponding ground truth summary using a suite of standard evaluation metrics, including Bleu score, chrF (character F1), Meteor, Sentence-BERT score, and Rouge.
-The results are noted in Table 3.
-We observe that across a majority of the metrics, PaliGemma-10B outperforms the other VLMs.
-Hence, for the purpose of the final tag extraction module in the next section, we shall use the summaries generated by the PaliGemma-10B model.
-Model Bleu chrF MT S-Bert Rouge
-PaliGemma-10B 0.1038 40.38 0.3188 77.65 0.2682
-Qwen-2.5 0.0821 42.30 0.3156 76.67 0.2443
-LLaVA-1.6 0.0292 18.10 0.1257 42.03 0.1117
-Table 3: Performance of different VLMs in generating tagged summaries across different metrics. Best results are highlighted. Here, MT: Meteor and S-Bert: Sentence-BERT.
-Tag generation
-We fine-tune the different VLMs – PaliGemma-10B, LLaVA-1.6, and Qwen-2.5 for generating the final tags.
-The fine-tuning prompt asks these VLMs to extract a set of tags from the summary of the meme generated by PaliGemma-10B and fed as a part of the prompt.
-The extracted output tags are compared with the ground truth tags of the meme and the loss is back propagated.
-The fine-tuning is done again on the training split noted in Table 2.
-Similarly, we test the performance of the tag generation model on the test split noted in Table 2.
-In the following we first discuss the baselines and the metrics used for evaluation.
-Baselines: We use three state-of-the-art baselines for evaluation.
-These include (i) Tag2Text (Huang et al.
-2024), (ii) RAM (Zhang et al.
-2024) and (iii) RAM++ (Huang et al.
-2023).
-Tag2Text, is a vision-language pre-training framework that incorporates image tagging to enhance the learning of visual-linguistic representations.
-Tag2Text extracts tags directly from associated text, enabling the model to learn an image tagger while simultaneously guiding the vision-language learning process.
-The RAM model performs image tagging by training on large-scale image-text pairs.
-An initial model is trained by jointly learning from image captions and tags, supervised respectively by the original textual data and the parsed tags.
-This is followed by a data refinement engine that generates new tags and removes noisy annotations.
-The model on this cleaned dataset is retrained and further fine-tuned on a smaller, high-quality set.
-RAM++ is an open-set image tagging model that effectively harnesses multi-granular textual supervision.
-RAM++ unifies individual tag-level and global text-level supervision within a single alignment framework.
-To further enhance tag understanding, it leverages LLMs to transform semantically narrow tag supervision into broader, descriptive tag supervision, thereby enriching the model’s grasp of visual concepts in open-set scenarios.
-Both RAM and RAM++ models can either use their own trained tags or the open-set vocabulary of tags during inference, resulting in two different variants.
-Metrics: We use multiple metrics to evaluate the similarity between the ground truth and the generated tags.
-These include the semantic similarity [^13] https://sbert.net/docs/sentence˙transformer/usage/semantic˙textual˙similarity.html, BERTScore and ConceptNet similarity [^14] https://conceptnet.io/.
-For each ground truth tag, we compute the semantic similarity (or BERT score or ConceptNet similarity) with all the generated tags and take the maximum value among these.
-Next, we average this maximum over all the ground truth tags.
-We average these averages over all the test data points.
-In addition, to capture better context, we also expand both the ground truth tag list as well as the generated tag list using the Google Search API and compare the expanded list using the same semantic similarity and BERTScore [^15] ConceptNet is primarily designed for obtaining one-word/short-phrase to one-word/short-phrase relationships.
-Hence, we have not used it to measure similarity between the expanded tag lists..
-Model Tag similarity Exp tag similarity
-$S$ BS ConceptNet $S$ BS
-PaliGemma-10B 63.9 93.55 51.40 54.5 95.44
-Qwen-2.5 59.1 92.65 45.23 48.2 94.77
-LLaVA-1.6 56.7 87.89 43.85 45.0 92.88
-RAM++ 40.6 90.36 19.43 24.9 92.48
-RAM++ Openset 27.8 82.22 8.54 12.8 91.57
-RAM 40.5 90.57 19.41 24.8 92.48
-RAM Openset 28.3 72.19 7.07 12.2 91.62
-Tag2Text 37.3 89.84 17.31 21.2 92.27
-Table 4: Performance of the tag generation module. Best results are highlighted. $S$: Semantic similarity, BS: BERTScore, Exp: Expanded.
+As a combined objective, we therefore propose a novel multitasking framework where we jointly train toxic tag generation and discriminative label classification.
+The main motivation of this framework is to leverage semantic supervision to guide discriminative classification where the embeddings are shaped into a representation space where the supervision provided by the generative task acts as a regularizer for the shared representation space.
+Our experimental observations (refer Section 7) reveal that supervising the model to generate explicit toxic tags stimulates the intermediate hidden representations of the model and boosts discriminative performance, thereby signalling that the model learns more detailed and meaningful representations in joint training.
+The overall loss is calculated as a combination of the generation and classification losses:
+$\mathcal{L}_{total}=\mathcal{L}_{gen}+\mathcal{L}_{cls}$ (6)
+where, $\mathcal{L}_{gen}$ represents the generative loss and $\mathcal{L}_{cls}$ represents the classification loss.
+This alignment makes the layer selection objective described in Section 6.1 more robust as a toxicity classifier.
+The details on the employed models and experimental setup are elaborated in Appendix F and G respectively.
+In contrast, the uni-tasking framework (see Table 4), considers only the classification loss $\mathcal{L}_{\text{cls}}$, since it is solely dedicated for toxicity detection.
+In the next section we detail out our emperical observations.
+7 Results
+In this section we assess the performance of StemTox in toxic meme detection on the test split of the ToxicTags dataset for both stages – I and II.
+We also present the results of tag generation performance achieved by StemTox.
+For our experiments, we use two open-source VLMs – PaliGemma and LLaVA (refer Appendix F and G for further details).
+Table 4 shows the experimental results.
+Toxic meme detection: StemTox with the multi-task learning objective consistently outperforms uni-task training across both models and stages, highlighting the benefit of joint learning for toxicity detection and tag generation, as summarized in Table 4.
+PaliGemma attains the highest macro F1 scores of 72.55 and 64.66 for stage I and II, respectively.
+This represents a notable improvement over the uni-task setting, which has macro F1 scores of 60.85 in stage I and 42.78 in stage II.
+A similar pattern is also observed for LLaVA, where StemTox in the multi-task framework achieves macro F1 scores of 67.24 and 57.99 for the two stages, outperforming the uni-task learning approch.
+Notably, StemTox outperforms the powerful GPT-4o model in a few-shot setup where the few-shot examples are selected based on the ground-truth tag similarity of the memes.
+To obtain these few-shot samples, we first compute the CLIP embeddings for each tag in the train set and also those that are associated with the query sample.
+Next, we find the maximum cosine similarity of each query tag embedding with the tag embeddings of an example meme from train set.
+For instance, given a test sample with tags $\{t_{1},t_{2},t_{3}\}$ and an example meme with tags $\{e_{1},e_{2},e_{3}\}$, we compute the cosine similarity between the CLIP embedding of tag $t_{1}$ and those of $e_{1}$, $e_{2}$, and $e_{3}$, and record the maximum similarity value.
+We repeat the same process for $t_{2}$ and $t_{3}$.
+The final tag similarity between the test and the example sample is then defined as the mean of the maximum similarity scores.
+Likewise StemTox also outperforms the setup where the last hidden layer is considered (LHL) for computing $\mathcal{L}_{cls}$ instead of the proposed entropy based scheme.
+Competing baselines and generalizability across datasets: To further ground StemTox, we compare it with the existing benchmarks and frameworks.
+In particular, we compare with two benchmark datasets (FHM and MAMI) and five baselines namely, PromptHate Cao et al.
+(2023b), ProCap Cao et al.
+(2023a), ModHate Cao et al.
+(2024b), Few-Shot Hee et al.
+(2024) and M2KE Lu et al.
+(2025).
+Both the FHM and MAMI datasets do not have tags that is necessary to train StemTox.
+For the training points of both these datasets we therefore infer the tags using a PaliGemma model that is fine-tuned on the ToxicTags dataset to predict tags.
+Next we use the gold labels and these inferred silver tags to train StemTox (PaliGemma backbone as it exhibits the best performance in Table 4) for each of the two datasets.
+Table 5 compares the results of the baselines with StemTox.
+Our approach shows substantial improvement over the other baselines across the FHM and MAMI datasets, achieving a macro F1 score of 78.11, 79.67 and accuracy of 78.4, 80.1 respectively.
+In contrast, the best-performing baseline (M2KE) respectively achieves a maximum accuracy and macro F1 score of 75.76 and 75.52 on FHM, and 75.85 and 75.71 on MAMI.
+Model FHM MAMI
+acc mF1 acc mF1
+PH 72.98 72.24 70.31 70.18
+PC 75.10 74.85 73.63 73.42
+MH 57.60 53.88 69.05 68.78
+FS 66.00 65.80 70.50 70.10
+MK 75.76 75.52 75.85 75.71
+StemTox 78.4 78.11 80.1 79.67
+Table 5: Results comparing our method with the previous works. Best results are highlighted. PH: PromptHate, PC: ProCap, MH: ModHate, FS: Few-Shot, MK: M2KE, mF1: Macro-F1, acc: Accuracy.
+model tag similarity
+chrF met $S$
+RAM++ 0.1942 0.058 0.406
+RAM++ (O) 0.094 0.012 0.279
+RAM 0.19 0.05 0.40
+RAM (P) 0.087 0.014 0.338
+RAM (O) 0.079 0.013 0.285
+Tag2Text 0.1631 0.041 0.372
+Tag2Text (P) 0.2102 0.087 0.332
+StemTox 0.4872 0.265 0.626
+Table 6: Performance of the tag generation module. Best results are highlighted. chrF: chr-F, met: Meteor, $S$: Semantic-Similarity, (O): open-set, (P): pre-trained using ToxicTags.
+Tag generation module: As stated earlier, Table 4 presents the tag generation results over ToxicTags dataset as well.
+The metrics used are chr-F, Meteor and Semantic-Similarity [^13] https://huggingface.co/tasks/sentence-similarity.
+For each ground-truth tag, we compute each of the above metrics with all the generated tags and take the maximum value among these.
+Next, we average this maximum over all the ground-truth tags and calculate mean of these averages over all the test data points.
+Among the two models used within the StemTox multi-tasking framework, PaliGemma consistently outperforms LLaVA in terms of chr-F, Meteor and Semantic-Similarity.
+Specifically, PaliGemma achieves scores of 0.4872 (chr-F), 0.265 (Meteor) and 0.626 (Semantic-Similarity), whereas LLaVA attains 0.2648, 0.107 and 0.447, respectively.
+These results indicate that PaliGemma is more effective than LLaVA for tag generation in the multi-tasking framework StemTox.
+Tag generation baselines: Since StemTox generate tags along with toxicity labels, it is imperative to compare its tag generation performance with state-of-the-art generators.
+To this purpose, we use three state-of-the-art baselines for evaluation namely Tag2Text Huang et al.
+(2024), RAM Zhang et al.
+(2024), and RAM++ Huang et al.
+(2023).
+To ensure fair evaluation, we further pre-trained the Tag2Text and RAM models on our ToxicTags dataset, allowing them to suitably adapt to domain-specific toxic vocabulary.
+Table 6 shows that StemTox by far outperforms all the other tag generation baselines in terms of all the three metrics – chr-F, Meteor and Semantic-Similarity.
+State-of-the-art models fail to capture the meme’s social context and their implicit meaning.
+They focus primarily on the visible objects in the image (refer to Table 7; also refer to Table 9 in Appendix for more examples).
 Image [Uncaptioned image] [Uncaptioned image]
-Ground truth trainers, hitler
+GT trainers, hitler
 911 9/11 twin towers impact,
@@
 9/11 truth movement
-Our approach hitler, trainers, nazi, just do it 911, twin towers
-RAM++ blue, man, shoe , smile building, city
-RAM++ Openset
+ST hitler, trainers, nazi 911, jenga, twin towers
+R++ blue, man, shoe , smile building, city
+R++(O)
 Athletic shoe, Close-up,
@@
 Close-up, Toy block
-RAM blue , man , shoe building, city
-RAM Openset
+R blue , man , shoe building, city
+R(P)
+R(O)
 Athletic shoe, Black-and-white,
@@
 Toy block
-Tag2Text
+T2T
 shoe, picture, uniform,
@@
 […]
-Table 5: Comparative examples of tags generated by different models. We provide more examples in Table 7 in the Appendix.
-Generation performance: The results of tag generation are presented in Table 4.
-The top three rows (our scheme) by far outperforms the state-of-the-art baselines in terms all the evaluation metrics.
-Among our models, PaliGemma-10B has the best tag generation performance.
-State-of-the-art models fail to capture the meme’s social context and their implicit meaning.
-They focus primarily on the visible objects in the image (refer Table 5; also refer Table 7 in Appendix for more examples).
-7 Toxic meme detection
-ToxicTags– In this section we assess the performance of toxic meme detection on the test split for both stages: Stage I and Stage II.
-Results portray the effectiveness of the tags in few-shot exemplar selection as we demonstrate below (see Table 6).
-Note that the details of the employed VLMs, embedding generation models, prompt and experimental setup are covered in Appendix FG and H respectively.
-Models Shots, Stage $_{\textrm{r}} I$$_{\textrm{im}}$ $_{\textrm{gt}} I$$_{\textrm{pt}}$ $_{\textrm{im}\oplus\textrm{gt}} I$$_{\textrm{im}\oplus\textrm{pt}}$
-2, I 51.95 56.94 56.83 58.19 59.66 61.03
-2, II 17.59 28.39 24.23 23.79 29.7 29.27
-4, I 46.5 62.18 58.51 60.44 62.25 63
-IDEFICS-3 4, II 17.58 39.27 35.52 35.45 41.6 41.4
-2, I 64.4 66.89 68 67.45 68.56 69.12
-2, II 44.29 50.68 53.28 51.41 55.6 54.47
-4, I 65.97 70.17 68.86 70.44 70.54 70.34
-GPT-4o 4, II 46.10 54.50 52.63 55.27 57.9 57.21
-IntMeme I – 45.85 II – 31.6
-Table 6: Macro-F1 scores for different few-shot example selection strategies evaluated on the ToxicTags dataset.
-Further, the final row shows the results for the latest and most competing baseline IntMeme (Hee and Lee 2025) evaluated on the ToxicTags dataset.
-Best: bold, second best: underline).
-Overall best result for each stage is highlighted.
-Shot selection: In the following we describe the various methods for few-shot selection.
+T2T(P)
+death, hitler, adolf hitler,
+holocaust, family,
 […]
+9/11, twin towers,
+911 9/11 twin towers
 […]
-[…]
-[…]
-[…]
-[…]
+Table 7: Examples of tags generated by different models.
+GT:Ground Truth, ST: StemTox, R: RAM, R++: RAM++, R(O): RAM open-set, R++(O): RAM++ open-set, T2T: Tag2Text, R(P): RAM pre-trained, T2T(P): Tag2Text pre-trained using ToxicTags.
+Please find more examples in Table 9 in the Appendix.
+8 Error analysis
+category PaliGemma LLaVA
+hateful 0.2288 $\pm$ 0.2088 0.3029 $\pm$ 0.1968
+dangerous 0.2423 $\pm$ 0.2043 0.3132 $\pm$ 0.1790
+normal 0.3221 $\pm$ 0.1670 0.3478 $\pm$ 0.1557
+offensive 0.4950 $\pm$ 0.1210 0.5559 $\pm$ 0.1065
+Table 8: Average uncertainty scores ($\pm$ standard deviation) across different toxicity categories for PaliGemma and LLaVA. Higher numbers signify higher uncertainty in the model’s classification.
+To understand the failures of the model we conduct an error analysis that helps to identify how these models struggle in classification; due to systematic weaknesses in contextual reasoning.
+We perform two complementary analyses to characterize these errors.
+(i) We identify the most frequent tags for which each model struggles to predict the correct classification labels.
+Key observations: For LLaVA, the majority of the misclassifications are concentrated around tags corresponding to death-related events (death, funeral, suicide, murder, heart attack), mental health and violence (depression, gun control, slavery), and religious or cultural entities (religion, Jesus, Bible), indicating a bias toward emotionally sensitive scenes.
+In the case of PaliGemma, higher error frequencies are observed for tags associated with humor and internet slang (e.g., comic, lol, oof size large, cursed), suggesting that the model is unable to effectively interpret implicit sarcasm and meme culture.
+(ii) To better understand where the models fail, we group all false positive samples according to their incorrectly predicted classes and perform an uncertainty-based error analysis.
+We quantify the model’s behavior using an uncertainty score derived from the Maximum Softmax Probability (MSP) Hendrycks and Gimpel (2017).
+This uncertainty score is defined as:
+$\mathcal{U}(x)=1-\max_{c}p_{\theta}(c\mid x)$ (7)
+where $x$ denotes a multimodal input sample, $c\in\{\textit{hateful},\textit{offensive},\textit{dangerous},\textit{normal}\}$ represents a toxicity class, and $p_{\theta}(c\mid x)$ is the model-predicted posterior probability given input $x$.
+Key observations: Table 8 shows that LLaVA consistently exhibits higher uncertainty for false positive predictions.
+Conversely, PaliGemma shows lower confidence margins.
+Both models display the highest uncertainty in the offensive category, indicating difficulty in learning accurate patterns, whereas interestingly the lowest uncertainty is observed in the hateful category.
 9 Conclusion
 This work makes several key contributions toward advancing research in toxic meme detection and multimodal content moderation.
-First, we curate a richly annotated dataset of 6,300 real-world meme-based posts through a two-stage labelling process.
+First, we curate a richly annotated dataset of 6,300 real-world meme-based posts through a two-stage labelling process which we call ToxicTags.
 Second, we enhance contextual understanding by introducing auxiliary metadata including meme titles and, most importantly, social tags.
-Third, we propose a generative AI-based tag generation module capable of producing socially relevant tags given an arbitrary input meme with no tag information in the first place.
-Finally, we evaluate the performance of state-of-the-art VLMs under few-shot prompting setups, establishing strong baselines for future research.
-Collectively, our contributions address key bottlenecks in toxic meme moderation and provide a foundation for building more accurate, context-aware, and socially responsible content moderation systems.
-Appendix A Limitations
+Third, we propose StemTox that detects toxicity as well as generates tags given an arbitrary input meme.
+Finally, we compare the performance of toxicity detection and tag generation on an array of benchmarks and baselines for generalizability.
+Collectively, our contributions address key bottlenecks in semantically aligned toxic meme moderation and thus provides a foundation for building more accurate, context-aware, and socially responsible systems.
+10 Limitations
 While this work offers several important contributions, it also has a few limitations.
@@
 Second, we do not examine the psychological and emotional impact of hateful memes on viewers.
-Exposure to such content may contribute to anxiety, depression, or desensitization—an important area that lies beyond the current scope.
+Exposure to such content may contribute to anxiety, depression, or desensitization – an important area that lies beyond the current scope.
 Addressing this limitation would require collaboration with psychologists, mental health experts, and affected communities.
-Third, we do not investigate the real-world consequences of online hate memes, particularly their potential to incite offline violence or criminal behavior.
+Third, we do not investigate the real-world consequences of online hate memes, particularly their potential to incite offline violence or criminal behaviour.
 Understanding this link between online toxicity and offline harm is a critical direction for future research.
-Appendix B Ethics statement
+Appendix A Ethics statement
 We strictly adhered to the policies of the social media platforms from which the dataset was curated.
@@
 Our work aims to support the broader effort to combat online hate while remaining mindful of its own limitations.
+Appendix B Prompt templates
+We display below the prompts used for Stage I and Stage II classifications.
+To generate the image caption, we first apply LLaMA image inpainting  Suvorov et al.
+(2021) to remove any OCR text from the image, and then prompt GPT to produce a caption.
+System prompt for stage I You are an AI assistant tasked with classifying memes.
+Consider the following definitions.
+$\bullet$ toxic: {definition of toxic}
+$\bullet$ normal: {definition of normal}
+Based on the above definitions, the input image, title, provided caption, list of tags and the extracted OCR text from the image delimited by three backticks classify the meme as toxic or normal.
+Example output for toxic meme: {toxic.}
+Example output for normal meme: {normal}
+System prompt for stage II You are an AI assistant tasked with classifying memes.
+Consider the following definitions.
+$\bullet$ hateful: {definition of hatful}
+$\bullet$ dangerous: {definition of dangerous}
+$\bullet$ offensive: {definition of offensive}
+$\bullet$ normal: {definition of normal}
+Based on the above definitions, the input image, title, provided caption, list of tags and the extracted OCR text from the image delimited by three backticks classify the meme as hateful, dangerous, offensive or normal.
+Example output for hatful meme: {hateful.}
+Example output for dangerous meme: {dangerous.}
+Example output for offensive meme: {offensive.}
+Example output for normal meme: {normal}
 Appendix C Platforms
-Data curation platform
-As stated earlier, we use imgflip.com as the curation platform for our data since it is centred on conversations using memes.
+C.1 Data curation platform
+As stated earlier, we use https://imgflip.com as the curation platform for our data since it is centred on conversations using memes.
 We specifically use the streams [^14] https://imgflip.com/streams feature of the platform to collect these memes.
 The selected streams and their descriptions are outlined as follows:
-  • Dark_Humour ($\sim$11k followers) [^17] https://imgflip.com/m/Dark˙humourWelcome to Dark_humour, Imgflip’s premier community for offensive humour.
+  • Dark_Humour ($\sim$11k followers) [^15] https://imgflip.com/m/Dark_humourWelcome to Dark_humour, Imgflip’s premier community for offensive humour.
 Stream mood: Relax liberals, it’s called dark humour.
-  • Memes_Overload ($\sim$9.5k followers) [^18] https://imgflip.com/m/MEMES˙OVERLOADHey there, welcome to MEMES_OVERLOAD, Imgflip’s 4th largest stream for memes!
+  • Memes_Overload ($\sim$9.5k followers) [^16] https://imgflip.com/m/MEMES_OVERLOADHey there, welcome to MEMES_OVERLOAD, Imgflip’s 4th largest stream for memes!
 We’re all here to have fun, so make sure to follow the rules, and keep on memeing on.
@@
 An example cropped post containing the corresponding title and tags is presented in Figure 4.
-Google Lens
-To provide contextual information from the real world, we provide two different augmentations to our tag generation module: (i) Google Search API based expansion of tags, and (ii) Google Lens based descriptions.
-While we have discussed in detail about tag expansion in the main content, here we outline the manual collection procedure and preprocessing applied for Google Lens context incorporation.
-Collection
-Figure 4: Example of manual collection of title and description of most relevant search obtained from Google Lens.
-Figure 4 shows an example meme with corresponding similar results obtained from Google Lens.
-We manually collect this information for each of the 6,300 memes by strictly asking the collectors to keep only the top-most matching result and to adhere to the following format for information storage: Title – Description.
-We believe that ours is the first of its kind utilisation of Google Lens for improving contextual knowledge.
-Preprocessing
-For pre-processing, we follow several steps to clean unwanted textual information from the information returned by Google Lens.
-First, we remove leading and trailing white spaces.
-Then we remove certain irrelevant URLs appended within the text, which could have generated ambiguity for the tag generation framework and further went on to clean some platform-specific formats such as subreddit references (e.g.
-r/, u/).
-Emojis are eliminated using a Unicode-based regex, and numerical metrics (e.g.
-2.5K likes, 3M views) & user mentions (e.g.
-@user) are also eliminated.
-Finally, attached metadata that do not contribute to meme semantics and non-English scripts (e.g.
-Chinese, Arabic, Cyrillic, and Devanagari) are also removed.
-We use ‘Unicode-range matching’ for targeting only English letters, numbers, and symbols while discarding other languages.
 Appendix D Details of annotation
-Figure 5: A step-by-step flowchart used by annotators to annotate any post.
 Figure 1 presents a brief pipeline of the employed annotation process over two stages.
 The instructions that we designed are given below, and some samples from those provided to annotators are present in Table 10.
-During the whole annotation process, annotators were asked to adhere to the definitions provided in the subsection D.
+During the whole annotation process, annotators were asked to adhere to the definitions provided in the subsection D.1.
   • Each annotator was provided with a separate PDF containing the meme, title, tags and the OCR extracted text from the meme.
@@
 All the annotations were performed by strictly adhering to the definition of hateful, dangerous and offensive labels.
-Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Image [Uncaptioned image] [Uncaptioned image]
 ground truth alabama, incest, baby chocolate, rivers, africa
-starving cannibal, neon lights,
-brothers, cannibalism
-Our approach alabama, abortion chocolate, rivers, africa
-starving, cannibal,
-cannibalism, cannibal
+StemTox alabama, problems, abortion, pregnancy chocolate, rivers, africa
 RAM++ anime, brush, girl, person, kiss, love
@@
 lay, man, mud, puddle, water
-barbecue grill, cook, dog, food,
-grill, person, man, meat
-RAM++ Openset Close-up Jeans Barbecue chicken, Filling station, Ribs
+RAM++ (O) Close-up Jeans
 RAM anime, girl, person, kiss
@@
 man, mud, puddle, water
-barbecue, barbecue grill, dog,
-food, grill , person, man, meat
-RAM Openset ’ ’ Ribs Ribs
+RAM (P)
+RAM (O) Ribs
 Tag2Text
-anime, girl, people,
+anime, girl, people, person
+User Specified
+mud, puddle, water, man
+User Specified
 […]
-mud, puddle, water,
-[…]
-grill, dog, person, man,
-[…]
-Table 7: Comparative examples of tags generated by different models.
+Table 9: Comparative examples of tags generated by different models. (O): open-set, (P): pre-trained using ToxicTags.
 Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
 Cannibal cafe menu If ykyk My friend sent me this
-Tags
-signs, advertising
+Tags signs, advertising
 hannibal lecter, cannibals,
@@
 the owl house, bee movie
-OCR extracted
-text
-imaflip.com ginger
-=== S
+OCR extracted text imaflip.com ginger === S
 CANNIBAL CAFE MENU BABY
-BACK RIBS FINGER SANDWHICH
+BACK RIBS FINGER SANDWICH
 OPEN FACE SANDWICH STU STEW
 TOE FU SCRAMBLED LEGS
-BAKED FRIARS MAC CHEESE
-( VEGAN )
+BAKED FRIARS MAC CHEESE (VEGAN)
 Consider John Frazzled
 FrazzleMyGimp PIZZA
-GUY : You’re total
-is \$ 26.34 ME : I can’t
-afford that PIZZA GUY :
-Well you’$l$ have to pay
-some other way ME :
-takes out wallet Wait I
-forgot I had 30 dollars
+GUY: Your total is \$26.34
+ME: I can’t afford that
+PIZZA GUY: Pay another way
 PORN DIRECTOR: Cut
@@
 this will confuse children
-But this is completely
-fine HOME TEETH 3
-Expert
-annotation
-hateful dangerous offensive normal
+But this is completely fine
+HOME TEETH 3
+Expert annotation hateful dangerous offensive normal
 Table 10: Four memes from the expert annotation samples provided to the annotators.
 [Uncaptioned image] [Uncaptioned image]
-hateful
-dangerous
+hateful dangerous
 [Uncaptioned image] [Uncaptioned image]
 offensive normal
-Table 9: Label wise word cloud of tags.
-Definitions
-hatefulReference(Kiela et al.
-2020) – A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease.
+Table 11: Label-wise word cloud of tags.
+D.1 Definitions
+hatefulReferenceKiela et al.
+(2020) – A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease.
 Attack is defined as violent or dehumanizing (comparing people to non-human things, e.g., animals) speech, statements of inferiority, and calls for exclusion or segregation.
@@
 dangerousReference [^18] https://www.dangerousspeech.org/dangerous-speech – A text, meme or speech which is not hateful but uses any form of expression that can increase the risk of its audience to condone or participate in violence against members of another group will be considered as dangerous.
-offensiveReferences(Roy et al. 2023; Mathew et al. 2022) – A text, meme or speech which is neither hateful, nor dangerous but uses abusive slurs or derogatory terms will be considered as offensive.
-toxicReference: PerspectiveAPI [^21] https://developers.perspectiveapi.com/s/about-the-api-model-cards?language=en˙US – A rude, disrespectful, or unreasonable comment that is likely to make you leave a discussion.
+offensiveReferencesRoy et al. (2023); Mathew et al. (2022) – A text, meme or speech which is neither hateful, nor dangerous but uses abusive slurs or derogatory terms will be considered as offensive.
+toxicReference: PerspectiveAPI [^19] https://developers.perspectiveapi.com/s/about-the-api-model-cards?language=en_US – A rude, disrespectful, or unreasonable comment that is likely to make you leave a discussion.
 normal – A meme which is not toxic and follows social norms.
-Safety guidelines
+D.2 Safety guidelines
 Primarily, we performed the following steps to keep our annotators mentally safe with such contents:
@@
 Time to go buy milk. Fast. But
-sonic . it’s YOUR baby !
+sonic, it’s YOUR baby!
 Ernie and Bert want to create
@@
 IT ONLY HURTS TIL YOU PASS OUT!
-MASKS SAVE LINES
+MASKS SAVE LIVES
+Ground truth
+annotation
+dangerous dangerous dangerous
 Table 12: Dangerous memes: portrayal of danger through anime/cartoon characters.
-tag pairs theme
-(‘hitler’, ‘nazi’), (‘hitler’, ‘jews’), (‘hitler’, ’ww2’), (‘hitler’, ‘holocaust’), (‘jews’, ‘nazi’), (‘concentration camp’, ‘nazi’), (‘hitler’, ‘stalin’) Historical events / World War II
-(‘9/11’, ‘twin tower’), (‘911’, ‘twin tower’), (‘9/11’, ’muslim’), (‘cow’, ‘muslim’), (‘911 9/11 twin tower impact’, ‘muslim’) 9/11 and Islamophobic references
-(‘jesus’, ‘satan’), (‘satan’, ‘the bible’), (‘jesus’, ‘the bible’) Religious contrast / satire
-(‘atomic bomb’, ‘hiroshima’), (‘hiroshima’, ‘ww2’), (‘hiroshima’, ‘nuke’) Nuclear warfare / World War II
-(‘alabama’, ‘incest’), (‘incest’, ‘sweet home alabama’), (‘alabama’, ‘dad’) Southern US stereotypes, taboo, humour
-(‘cannibal’, ‘cannibalism’), (‘fresh’, ‘meat’), (‘cannibalism’, ‘meat’), (‘human’, ‘meat’) Creepy / disturbing themes
-(‘mass shooting’, ‘school shooting’), (‘gun’, ‘school shooting’), (‘gun’, ‘school’), (‘gun’, ‘usa’), (‘mass shooting’, ‘usa’), (‘mass shooting’, ‘america’) US gun violence
-(‘baby’, ‘dead’), (‘baby’, ‘cannibalism’) Child-related violence / abortion
-(‘africa’, ‘water’), (‘africa’, ‘hungry’), (‘africa’, ‘starve’) Humanitarian crises in Africa
-(‘black people’, ‘racist’), (‘black’, ‘black life matter’), (‘black’, ‘white’), (‘angry black guy’, ‘lame’) Racism
-(‘elmo’, ‘sesame street’), (‘big bird’, ‘sesame street’), (‘ernie’, ‘sesame street’), (‘mayor’, ‘serial killer’), (‘free candy van’, ‘sonic the hedgehog’) Cartoon / anime characters in dark contexts
-(‘world war 3’, ‘ww3’), (‘ukraine’, ‘ww3’), (‘gaza’, ‘israel’), (‘monster’, ‘ww3’) Global conflict / war escalation
-Table 11: Top co-occurring tag pairs and associated semantic themes.
-Related tags: To understand the semantic themes associated with frequently occurring tag pairs, we conduct a co-occurrence analysis.
-Prior to the analysis, all tags are lemmatised to ensure consistency and improve result accuracy.
-Our primary objective is to identify how often specific tags co-occur and what thematic contexts they represent within the meme content.
-The key observations are as follows.
-Some of the top co-occurring pairs and their themes are listed in Table 11.
 Word clusters of tags: Figure 11 presents the word cloud of tags for hateful, dangerous, offensive and normal memes.
@@
 For offensive and normal, even though they have similar words, the presence of terms like cursed, nsfw and usage of comic character terms like sesame, street significantly separates them.
-Figure 6: Prompts used for Stage I and Stage II classification with few-shot exemplars.
 Appendix F Employed models
-IDEFICS-3: The latest version of VLM released by HuggingFace team, this is among the few open-source models that are trained with few-shot training objective.
-We use the instruction fine-tuned version for conducting our experiments – HuggingFaceM4/idefics-9b-instruct.
-LLaVA-1.6: An enhanced version of LLaVA (Liu et al.
-2023b) and LLaVA-1.5 (Liu et al.
-2023a), LLaVA-1.6 has shown significant improvement over its prior models.
-We use llava-hf/llava-v1.6-mistral-7b-hf checkpoint for running our experiments.
 PaliGemma-2: A successor of PaliGemma, PaliGemma-2 presents a series of models of varying size with further improved capabilities due to the incorporation of Gemma-2 LLM instead of Gemma.
 We use google/paligemma2-10b-pt-224 version from HuggingFace in this work.
-Qwen-2.5 VL: Successor of Qwen VL and Qwen-2 VL, this latest iteration introduces streamlined and efficient vision encoder and dynamic resolution for video understanding.
-We use the instruction-tuned checkpoint from HuggingFace – Qwen/Qwen2.5-VL-7B-Instruct in our experiments.
-GPT-4o: One of the first models by OpenAI team to have multimodal capability, GPT-4o (OpenAI 2024) has proved its wide applicability across multiple domains [^23] https://learn.microsoft.com/en-us/azure/ai-services/openai/.
+LLaVA-1.5: LLaVA-1.5 Liu et al. (2023) has shown significant improvement over its prior models. We use llava-hf/llava-1.5-7b-hf checkpoint for running our experiments.
+GPT-4o: One of the first models by OpenAI team to have multimodal capability, GPT-4o OpenAI (2024) has proved its wide applicability across multiple domains [^21] https://learn.microsoft.com/en-us/azure/ai-services/openai/.
 License agreement: We agreed to the terms of usage of all employed VLMs before using them in our work.
-Appendix G Employed prompts
-Detailed list of prompts for the corresponding variants for Stage-I and Stage-II are provided 6.
-Appendix H Experimental setup
-(i) Experimental setup for tag generation module:
-(a) For generating caption using the LaMa inpainted image, we set max new tokens to 30.
-(b) For generating tagged summary we have used maximum number of tokens to 150.
-(c) For finetuned summary generator and tag generation module, all the models are finetuned using Trainer API, where we set batch size as 1, gradient accumulation steps set to 1 and a learning rate of 2e-5.
-We use AdamW Torch optimizer with a $\beta_{2}$ value of 0.999 and a weight decay of 1e-6.
-Transformer modules including q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj and down_proj, with a rank value of 8 and task type set to causal language modeling, are finetuned using LoRA-based parameter-efficient fine-tuning.
-We use 4-bit quantization with bfloat16 compute precision through the BitsAndBytesConfig for efficient memory usage and faster training.
-To stabilize the initial learning phase, we set a warmup step of two and the number of training epochs to three.
-To control the maximum sequence length, we set max length to 512, while truncation and padding are also applied.
-(ii) Few-shot prompting: To control the randomness of the generated output, we set the temperature to 0.001 for all models used in Stage I and II.
-The max new tokens was set to 30 and all other parameters were kept as default.
+Appendix G Experimental setup
+We train both, uni- and multi-tasking framework using a consistent experimental configuration to ensure a fair comparison.
+For multi-tasking, the training objective consists of two components: a weighted cross-entropy loss for the classification task and the default autoregressive generation loss for toxic tag generation.
+For uni-tasking there is no generation loss.
+For computing class weights, we follow the weighting scheme described in King and Zeng (2001).
+All input sequences are truncated or padded to a fixed maximum sequence length of 2048 tokens, and the random seed is fixed to 42 to ensure reproducibility.
+Training is performed for three epochs with a per-device batch size of 1 and a gradient accumulation step of 4, resulting in an effective batch size of 4.
+We use the AdamW optimizer with a learning rate of $2\times 10^{-5}$, weight decay of $1\times 10^{-6}$, and $\beta_{2}$ set to 0.999, along with a warm-up of two steps.
+For memory-efficient training, the model is loaded using 8-bit quantization via the BitsAndBytes configuration, with computations carried out in FP16 precision.
+For the multi-task Learning framework, we apply LoRA-based parameter-efficient fine-tuning with rank 8, targeting the following modules: {q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj}.
+In contrast, for the uni-task learning framework, the model backbone is fully frozen during training.
+Extracted hidden representations are normalized using the default normalization layers provided by the backbone model at the time of forward propagation.
+A sequence compression module is used to extract features into a fixed-dimensional embedding, which is then passed through a label classification head to predict the final class labels.
+For the ToxicTags dataset, we fix the number of classification classes to four and for the FHM and MAMI datasets, we fix the number of classes to two.
+For result calculation and analysis of stage I, we merge the hateful, dangerous, and offensive categories into a single toxic class.
+This aggregation is the reverse of the two-stage annotation strategy followed during dataset construction.